### v2

# Exploiting Observation Bias to Improve Matrix Completion

Abstract
We consider a variant of matrix completion where entries are revealed in a biased manner, adopting a model akin to that introduced by Ma and Chen (2019). Instead of treating this observation bias as a disadvantage, as is typically the case, the goal is to exploit the shared information between the bias and the outcome of interest to improve predictions. Towards this, we consider a natural model where the observation pattern and outcome of interest are driven by the same set of underlying latent or unobserved factors. This leads to a two stage matrix completion algorithm: first, recover (distances between) the latent factors by utilizing matrix completion for the fully observed noisy binary matrix corresponding to the observation pattern; second, utilize the recovered latent factors as features and sparsely observed noisy outcomes as labels to perform non-parametric supervised learning. The finite-sample error rates analysis suggests that, ignoring logarithmic factors, this approach is competitive with the corresponding supervised learning parametric rates. This implies the two-stage method has performance that is comparable to having access to the unobserved latent factors through exploiting the shared information between the bias and outcomes. Through empirical evaluation using a real-world dataset, we find that with this two-stage algorithm, the estimates have 30x smaller mean squared error compared to traditional matrix completion methods, suggesting the utility of the model and the method proposed in this work.
1 Introduction
Matrix Completion. Matrix completion is a core learning task in many modern applications such as clinical trials Hazewinkel et al. (2022), recommender systems Steck (2010), and public policy evaluation Abadie et al. (2010), to name a few. It is concerned with estimating a matrix from partial and noisy observations of its entries. More specifically, the task consists in estimating an $n\times n$ matrix of interest $\mathbf{X}=(x_{ij})_{(i,j)\in[n]\times[n]}$, where only a noisy subset of its entries is observed. Such subset of entries may be characterized by a mask or a so-called observation matrix $\mathbf{A}=(a_{ij})_{i,j\in[n]\times[n]}$ taking values in $\{0,1\}$, i.e., if $a_{ij}=1$, then that a noisy version $y_{ij}$ of $x_{ij}$ is observed, otherwise the entry $(i,j)$ is considered missing.
Missing Completely At Random (MCAR). Due to its relevance, the matrix completion problem has been extensively studied in the recent years, yet most existing work Keshavan et al. (2010, 2009); Candes and Recht (2012); Candès and Tao (2010); Recht (2011); Negahban and Wainwright (2009); Davenport et al. (2014); Chen and Wainwright (2015); Chatterjee (2015); Xu (2017); Abbe et al. (2020) only focused on the so-called missing completely at random (MCAR) model. In this model, all entries are revealed independently with $a_{ij}\sim\mathsf{Ber}(p)$ for some probability parameter $p$. These works propose consistent estimators provided certain conditions on $p$ are satisfied and $\mathbf{X}$ is of low-rank $d\ll n$ with further structural assumptions. Notably, Keshavan et al. (2010) showed that spectral methods with refinement give consistent estimators provided that $p=\Omega(n^{-1}\log(n))$, and Borgs et al. (2022) proposed an iterative approach based on collaborative filtering which is also consistent in a stronger sense if $p=\Omega(n^{-1+\kappa})$ for some $\kappa>0$.
Missing Not At Random (MNAR). This work departs from the MCAR framework and instead concerns the practical and more general setting where entries may be missing not at random (MNAR). In this model, entries are revealed independently with $a_{ij}\sim\mathsf{Ber}(p_{ij})$ where $p_{ij}$ may even be zero. Indeed, in many real world applications, as is the case in clinical trials Steck (2010) or recommender systems Hazewinkel et al. (2022), the pattern of missing entries may carry some bias that may explain the value of the missing entries. For example, consider the case of restaurant recommendations where $\mathbf{X}$ would correspond to a rating matrix. We would not expect a vegan to visit and rate a steakhouse. Similarly, a recommendation algorithm may not recommend steakhouse to a vegan potentially based on their prior interactions.
Indeed, the presence of such bias in observations is an example of the classical confounding. The factors driving the reasons for not observing an outcome are also influencing the outcome itself. In the setting of matrix completion, however, such confounders are unobserved. This is precisely what has driven recent interest in matrix completion in the MNAR data model. We recall some of the recent and relevant prior works on it next.
Relevant Prior Work on MNAR. We briefly recall recent works from Ma and Chen (2019); Agarwal et al. (2023); Bhattacharya and Chatterjee (2022). Detailed relevant discussions on it can be found therein. We note that Agarwal et al. (2023) is closet to the setup considered here and we shall point to differences between them subsequently.
In Ma and Chen (2019), the propensity matrix $\mathbf{P}=\mathbb{E}[\mathbf{A}]$ is assumed to have low nuclear norm and is estimated directly. It is then used to recover the outcome matrix. Therefore, it requires $p_{ij}$ to be uniformly bounded away from 00, while our work does not impose such a restriction. We also refer an interested reader to other works taking similar approaches Yang et al. (2021); Liang et al. (2016); Schnabel et al. (2016); Wang et al. (2020).
In Bhattacharya and Chatterjee (2022), the authors introduce the self-masking MNAR model, where the observation pattern is a scalar function of the outcomes themselves. Their model can be viewed as a special case of the model introduced in this work. Additionally, our method provides sharper error rates compared to their work. We also refer an interested reader to other works taking similar approach Sportisse et al. (2020b, a).
In Agarwal et al. (2023), the authors consider the model closest to this work where the observation pattern and outcomes are determined by the same set of latent factors. The outcomes have low-rank representation with respect to the latent factors while the observation pattern has generic dependence on the latent factors. However, the dependence of the observation pattern is implicitly constrained by the requirement imposed by their Synthetic Nearest Neighbors algorithm, which requires a certain block pattern of observations. The functional form of this observation pattern is absent beyond such a constraint, thus failing to indicate what functional form would actually induce such patterns.
In comparison, as formally described below, we require that observation patterns have low-rank representation with respect to latent factors while outcomes have a Lipschitz function (with low-rank spectrum) dependence on the latent factor. In a sense, this work can be viewed as a way to formalize both the observation pattern and outcome pattern functionally in the model introduced in Agarwal et al. (2023). Indeed, this allows us to compare the performance of our method with a natural lower bound that is suggested by parametric rate of supervised learning. Such was absent in Agarwal et al. (2023) due to lack of formalizing observation pattern structure.
Contributions.
Below, we summarize the key contributions of our work:
  • (a)
    First, we introduce a novel model in the MNAR framework that is richer, more practical, and more expressive than existing ones. We consider that there is shared information between the missingness pattern and outcomes of interest. More specifically, the probability of observation is $p_{ij}\propto n^{-\beta}(1+\mathbf{u}_{i}^{\top}\mathbf{v}_{j})$ for some $\beta>0$ where $\mathbf{u}_{i},\mathbf{v}_{j}$ are $d$-dimensional latent vectors. The outcome of interest $x_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})$ where $f$ is a Lipschitz function that admits a low-rank representation of order $d$. We refer the reader to section 2 for the precise presentation of our model.
  • (b)
    Second, we propose a novel two-stage algorithm, referred to as Mask Nearest Neighbors (MNN): (i) in the first stage, we build estimates of the distances between the latent vectors, then use these estimates to perform clustering and denoising of the observations; (ii) in the second stage, we use the revealed entries to impute the missing ones. We establish that MNN achieves, under our shared information model, a finite-sample entry-wise performance guarantee of order $\widetilde{O}(n^{-(2-\beta)/(2d)})$ (see Theorem 4.1).
  • (c)
    Third, the prediction error rates achieved by MNN is competitive with parametric rates for corresponding supervised learning methods despite not knowing the latent vectors, and dealing with an unconventional observation model. More precisely, MNN achieves an error rate of order $\widetilde{O}(N^{-1/(2d)})$ which is comparable to the optimal parametric rate for the corresponding supervised learning task known to be of order $\widetilde{O}(N^{-1/(2+2d)})$ where $N$ denotes the number of observed entries (see Lemma 4.1 and discussion in Section 4.2).
  • (d)
    Finally, we establish empirically the performance gains promised by our theoretical results in the real-world setting. Specifically, we utilize more than a million interaction data points of users on the Glance platform – a smart lock-screen that aims to personalize user experience through recommending dynamic lock screens (also called glances). We find that MNN can improve the mean-squared error by 30x compared to a standard matrix completion method, as can be seen in Table 2. We also report empirical performance using a synthetic dataset to discuss nuanced properties of MNN that are not necessarily captured by theoretical results.
Notation.
We recall $\mathcal{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$. We write $[n]$ to mean $\{1,\dots,n\}$. For a given parameter $p\in[0,1]$, $\mathsf{Ber}(p)$ denotes the Bernoulli distribution with parameter $p$. For any $a,b\in\mathbb{R}$, we write $a\wedge b=\min(a,b)$ and $a\vee b=\max(a,b)$. For any given vector $\mathbf{x}\in\mathbb{R}^{d}$, we denote its euclidean by $\lVert\mathbf{x}\rVert$. For a given $m\times n$ matrix $\mathbf{M}$, for any subsets $\mathcal{I}\subseteq[n]$, $\mathcal{J}\subseteq[m]$, we denote $\mathbf{M}_{\mathcal{I},\mathcal{J}}$ the sub-matrix of $\mathbf{M}$ constructed by removing the rows $[n]\backslash\mathcal{I}$ and columns $[m]\backslash\mathcal{J}$. We further use $\mathbf{M}_{i,:}$ (resp. $\mathbf{M}_{:,j}$) to denote the $i^{\text{th}}$ row (resp. $j^{\text{th}}$ column) of $\mathbf{M}$. We denote the max norm of $\mathbf{M}$ by $\|\mathbf{M}\|_{\max}=\max_{(i,j)\in[m]\times[n]}|\mathbf{M}_{i,j}|$, its operator norm by $\|\mathbf{M}\|_{\textup{op}}$, its Frobenius norm by $\|\mathbf{M}\|_{\textup{F}}$, and its two-to-infinity norm by $\|\mathbf{M}\|_{2\to\infty}=\max_{i\in[m]}\|\mathbf{M}_{i,:}\|$. The pseudo-inverse of the matrix $\bm{M}$ is denoted by $\bm{M}^{\dagger}$.
2 Problem Description
2.1 Model and learning task
Latent factor model.
We posit that users (rows) and items (columns) can be fully described by $d$-dimensional unit vectors where $d\ll n$. More specifically, each user $i$ (resp. item $j$) is associated with a unit vector $\mathbf{u}_{i}\in\mathcal{S}^{d-1}$ (resp. $\mathbf{v}_{j}\in\mathcal{S}^{d-1}$), which we assume to be distributed independently and uniformly at random from the unit sphere. We will refer to $(\mathbf{u}_{i})_{i\in[n]}$ (resp. $(\mathbf{v}_{j})_{j\in[n]}$) as the user (resp. item) latent factors. Such latent factors determine both the probability of observation as well as the outcome of interest.
Outcomes.
The outcomes of interest may be interpreted as affinity scores between user-item pairs. In our model, we denote the outcome of each user-item pair $(i,j)\in[n]\times[n]$ by $x_{ij}$, and consider that it may be expressed as a sum of $d$ separable functions:
$x_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})=\sum_{\ell=1}^{d}\theta_{\ell}(\mathbf{u}_{i})\phi_{\ell}(\mathbf{v}_{j}),$ (1)
where the mappings $\theta(\cdot)=(\theta_{1}(\cdot),\dots,\theta_{d}(\cdot))$ and $\phi(\cdot)=(\phi_{1}(\cdot),\dots,\phi_{d}(\cdot))$ are defined from $\mathcal{S}^{d-1}$ to $[-B,B]^{d}$, and are assumed to be $L$-lipschitz. As such, the outcomes, despite forming a low-rank matrix of rank $d$, are not simply bi-linear forms of the user and item latent factors as is traditionally the case in matrix completion.
Observation model.
Our setup differs from the classical matrix completion literature in that entries are not randomly revealed with uniform probability. Instead, the observation probability of a given user-item pair $(i,j)$ is defined as:
$p_{ij}=\frac{\rho_{n}}{2}(\mathbf{u}_{i}^{\top}\mathbf{v}_{j}+1),$ (2)
where $\rho_{n}\in(0,1]$ is a sparsity factor that may vary with $n$. This means that the observation pattern induced is very closely related to the random dot product graph construction Athreya et al. (2018). Considering the infamous selection bias issue in recommender systems, which is well-displayed in Ma and Chen (2019), this is a more realistic assumption. The observation indicator is then $a_{ij}\sim\mathsf{Ber}(p_{ij})$ and independent across user-item pairs. Having $a_{ij}=1$ means that the entry $(i,j)$ is revealed, and we observe a noise-corrupted outcome $y_{ij}$. More precisely, we have
$y_{ij}=\begin{cases}x_{ij}+\xi_{ij}&\textrm{if }a_{ij}=1,\\ \star&\textrm{otherwise}\end{cases}$ (3)
where we write $\star$ to mean that an entry is not revealed. $(\xi_{ij})_{i,j\in[n]}$ are i.i.d. copies of a zero-mean and $\sigma^{2}$-sub-gaussian random variable.
Objective.
Our objective is that of prediction error minimization: Given the observation pattern $(a_{ij})_{i,j\in[n]}$, and observations $\mathcal{D}=\{y_{ij}:a_{ij}=1,i,j\in[n]\}$, we aim at predicting the outcomes $(x_{ij})_{i,j\in[n]}$ with finite-sample error rates at a user-item level. More specifically, we wish to construct estimates $(\hat{x}_{ij})_{i,j\in[n]}$ with small prediction error $|\hat{x}_{ij}-x_{ij}|$, for all $i,j\in[n]$, in high probability.
In the subsequent sections, we will often prefer matrix notations. We denote $\mathbf{U}^{\top}=\begin{bmatrix}\mathbf{u}_{1}&\cdots&\mathbf{u}_{n}\end{bmatrix}$, $\mathbf{V}^{\top}=\begin{bmatrix}\mathbf{v}_{1}&\cdots&\mathbf{v}_{n}\end{bmatrix}$ in $\mathbb{R}^{d\times n}$, $\mathbf{X}=(x_{ij})_{i,j\in[n]}$, $\mathbf{P}=(p_{ij})_{i,j\in[n]}$, $\mathbf{A}=(a_{ij})_{i,j\in[n]}$ and $\mathbf{Y}=(y_{ij})_{i,j\in[n]}$ in $\mathbb{R}^{n\times n}$.
2.2 Rationale behind the proposed model
Presence of confounding.
The model we propose captures a phenomenon that often occurs in practice. As an example, consider the case of the Glance smart lock-screens, where the rows of $\mathbf{X}$ correspond to users, the columns correspond to individual glances, and an entry $x_{ij}$ corresponds to the time user $i$ spends interacting with glance $j$. We note here that a user not interested in sports is unlikely to spend time viewing a glance recounting a recent basketball game. Moreover, based on the behavior of this user, the recommendation algorithm is unlikely to recommend glances about sports. Hence, information about the users’ preferences can be extracted both from their engagement time and which lock screens they are shown. This shared information corresponds to common latent features that affect both the observation pattern and the outcome in our model.
Low-rank representation.
In view of the above discussion, we may hypothesize that both the observation probabilities and outcomes of interest may be expressed as $p_{ij}=g(\mathbf{z}_{i},\mathbf{w}_{j})$ and $x_{ij}=h(\mathbf{z}_{i},\mathbf{w}_{j})$ for all $i,j\in[n]$, where $g$ and $h$ are both “nice" functions of shared latent factors $\mathbf{z}_{i},\mathbf{w}_{j}$ in $\mathbb{R}^{d}$. Here, by “nice" functions, we could mean for instance that $g$ and $h$ are bounded and Lipschitz. This is a reasonable assumption under which we may use a classical spectral decomposition Conway (2019). Below, we present a variant of this decomposition which we borrow from Shah et al. (2020) (see their Theorem 1).
Theorem 2.1.
Let $h:\mathcal{S}^{d-1}\times\mathcal{S}^{d-1}\mapsto\mathbb{R}$ be bounded and $L$-Lipschitz with respect to the $1$-product metric in $\mathcal{S}^{d-1}\times\mathcal{S}^{d-1}$. Then, there exists a non-increasing sequence $(\sigma_{k})_{k\geq 1}$ with $\sum_{k=1}^{\infty}\sigma_{k}^{2}<\infty$, and orthogonal families of functions $\{\alpha_{k}\in L^{2}(\mathcal{S}^{d-1}):k\in\mathbb{N}^{\star}\}$, and $\{\gamma_{k}\in L^{2}(\mathcal{S}^{d-1}):k\in\mathbb{N}^{\star}\}$ such that, for all $\mathbf{z},\mathbf{w}\in\mathcal{S}^{d-1}$,
$h(\mathbf{z},\mathbf{w})=\sum_{k=1}^{\infty}\sigma_{k}\alpha_{k}(\mathbf{z})\gamma_{k}(\mathbf{w}).$ (4)
Now using Theorem 2.1, we may argue that both $g$ and $h$ admit a low-rank representation, say of rank $d$ for simplicity, where we write $p_{ij}=g(\mathbf{z}_{i},\mathbf{w}_{j})=\sum_{\ell=1}^{d}u_{\ell}(\mathbf{z}_{i})v_{\ell}(\mathbf{w}_{j})$, and $x_{ij}=h(\mathbf{z}_{i},\mathbf{w}_{j})=\sum_{\ell=1}^{d}\alpha_{\ell}(\mathbf{z}_{i})\gamma_{\ell}(\mathbf{w}_{j})$ for all $i,j\in[n]$ with $u(\cdot),v(\cdot),\alpha(\cdot)$, and $\gamma(\cdot)$ are mappings from $\mathcal{S}^{d-1}$ to $\mathbb{R}^{d}$. Thus, if additionally $u(\cdot)$ and $v(\cdot)$ are invertible, then we may define $\theta=\alpha\circ u^{-1}$ and $\phi=\gamma\circ v^{-1}$ and obtain $p_{ij}\propto\mathbf{u}_{i}^{\top}\mathbf{v}_{j}$ and $x_{i,j}=f(\mathbf{u}_{i},\mathbf{v}_{j})=\sum_{\ell=1}^{d}\theta_{\ell}(\mathbf{u}_{i})\phi_{\ell}(\mathbf{v}_{j})$, where $\mathbf{u}_{i}=u(\mathbf{z}_{i})$ and $\mathbf{v}_{j}=v(\mathbf{w}_{j})$ for all $i,j\in[n]$. With this rough argument, we recover a low-rank representation similar to the one described in our model.
3 The Mask Nearest Neighbor (MNN) Algorithm
We propose the Mask Nearest Neighbor (MNN) algorithm that runs in two stages: (i) The first stage consists in reducing the noise in the observed outcomes by clustering users and items that are sufficiently close to each other, then averaging over these clusters whenever there are enough entries observed within a cluster. This averaging leads to a new matrix with a reduced size and still missing entries; (ii) The second stage consists in imputing the missing entries of the reduced matrix, then extrapolating from it to predict all outcomes.
3.1 Clustering & Denoising
Distance estimation.
The first step of the algorithm is to estimate the user-user and item-item distances in the latent space, denoted $d_{ij}^{\mathbf{u}}=\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}$, and $d_{ij}^{\mathbf{v}}=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|_{2}$, for all $i,j\in[n]$, respectively. To that end, we start by constructing the centered matrix of observations $\widetilde{\mathbf{A}}=\mathbf{A}-(\rho_{n}/2)\bf 1\bf 1^{\top}$, which we then project onto the space of rank $d$ matrices. More precisely, given the singular value decomposition $\widetilde{\mathbf{A}}=\sum_{i=1}^{n}\sigma_{i}\mathbf{q}_{i}\mathbf{w}_{i}^{\top}$ we construct an estimate of $\widetilde{\mathbf{P}}=\mathbb{E}[\widetilde{\mathbf{A}}]$ as follows:
$\widehat{\widetilde{\mathbf{P}}}=\sum_{i=1}^{d}\sigma_{i}\mathbf{q}_{i}\mathbf{w}_{i}^{\top},$ (5)
where $\sigma_{1}\geq\dots\geq\sigma_{n}$ are the singular values of $\widetilde{\mathbf{A}}$, and $\mathbf{q}_{i}$ and $\mathbf{w}_{i}$ are the corresponding left and right singular vectors. Then, for users (resp. items) $i$ and $j$ we can then define the distance estimate
$\hat{d}^{\mathbf{u}}_{ij}$ $=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}\left\lVert\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widehat{\widetilde{\mathbf{P}}}_{j,:}\right\rVert_{2},$
$\hat{d}^{\mathbf{v}}_{ij}$ $=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}\left\lVert\widehat{\widetilde{\mathbf{P}}}_{:,i}-\widehat{\widetilde{\mathbf{P}}}_{:,j}\right\rVert_{2}.$ (6)
Clustering.
We group the users and items by first selecting maximal subsets that are sufficiently separated. Concretely, we initialize the central users, denoted $\mathcal{U}^{\star}\subseteq[n]$, as the empty set. Then, we repeatedly add an arbitrary user to the set if its estimated distance from any other central user is at least $6\varepsilon_{n}$, where $\varepsilon_{n}$ is a hyperparameter we defer quantifying. $\mathcal{V}^{\star}$ is defined analogously.
By assigning each non-central user to its closest central counterpart in estimated distance, a partition of the users $\bigcup_{i\in\mathcal{U}^{\star}}\mathcal{U}_{i}=[n]$ is formed. Similarly construct the item partition $\bigcup_{j\in\mathcal{V}^{\star}}\mathcal{V}_{j}$. Note that we index the partitions by the sets $\mathcal{U}^{\star},\mathcal{V}^{\star}$ to avoid confusion from multiple indexing systems.
The observations can then be partitioned as $\mathcal{D}=\bigcup_{i\in\mathcal{U}^{\star}}\bigcup_{j\in\mathcal{V}^{\star}}\{y_{k\ell}:(k,\ell)\in\Omega_{ij}\}$, where $\Omega_{ij}=\{(k,\ell):a_{k\ell}=1,k\in\mathcal{U}_{i},\ell\in\mathcal{V}_{j}\}$ is the set of observed outcomes between users in cluster $i$ and items in cluster $j$.
Denoising.
At this stage, we are ready to define the ground truth clustered outcome matrix $\mathbf{H}=(h_{ij})_{i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}}$, where
$h_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j}),\quad i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}.$ (7)
We again emphasize that the rows and columns of $\mathbf{H}$ are indexed by central users $\mathcal{U}^{\star}$ and central items $\mathcal{V}^{\star}$ for simplicity. Using data, we form the unfilled clustered outcome matrix $\overline{\mathbf{H}}\in\mathbb{R}^{|\mathcal{U}^{\star}|\times|\mathcal{V}^{\star}|}$, where for all $(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$,
$\bar{h}_{ij}=\begin{cases}\frac{1}{|\Omega_{ij}|}\sum_{(k,l)\in\Omega_{ij}}y_{kl}&\quad\text{if $|\Omega_{ij}|\geq N_{n}$;}\\ \star&\quad\text{otherwise.}\end{cases}$ (8)
$N_{n}$ is a hyperparameter that governs whether there are “enough” observations available to produce a sufficiently accurate estimate of $h_{ij}$ via simple averaging. The choice of this hyperparameter will be specified later.
3.2 Imputation & Outcome Prediction
Imputation.
We will impute the missing entries of $\overline{\mathbf{H}}$ using the available entries. We will define the imputed clustered outcome matrix $\widehat{\mathbf{H}}$, with
$\hat{h}_{ij}=\begin{cases}\bar{h}_{ij}&\quad\text{if $\bar{h}_{ij}\neq\star$};\\ \texttt{Impute}(\overline{\mathbf{H}},i,j)&\quad\text{otherwise.}\end{cases}$ (9)
The subroutine $\texttt{Impute}(\overline{\mathbf{H}},i,j)$ proceeds as follows:
  • (i)
    First, we search over the sets $\mathcal{U}^{\star}$ and $\mathcal{V}^{\star}$ to find subsets of $d$ distinct central users $\mathcal{I}$ and $d$ distinct central items $\mathcal{J}$ such that $\sigma_{r}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})$ is maximized and
    $\!\!\!\!\!\!\!\!\!\!\!\!\!\!\forall(i^{\prime},j^{\prime})\in\mathcal{I}\times\mathcal{J},\quad\bar{h}_{i,j^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime},j^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime},j}\neq\star.$ (10)
  • (ii)
    Then, we define
    $\!\!\!\!\!\!\!\!\!\!\!\!\texttt{Impute}(\overline{\mathbf{H}},i,j)\!=\!\!\begin{cases}\!\overline{\mathbf{H}}_{i,\mathcal{J}}\!\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\!\right)^{\!\dagger}\!\overline{\mathbf{H}}_{\mathcal{I},j}&\!\!\!\!\text{if }\mathcal{I},\mathcal{J}\text{ are found};\\ \star&\!\!\!\!\text{otherwise.}\end{cases}$ (11)
Outcome Prediction.
Once $\widehat{\mathbf{H}}$ is constructed, predicting outcomes is simple: For any $i,j\in[n]$, there exists $k\in\mathcal{U}^{\star}$ and $\ell\in\mathcal{V}^{\star}$ such that $i\in\mathcal{U}_{k}$ and $j\in\mathcal{V}_{\ell}$, we then set $\hat{x}_{ij}=\hat{h}_{kl}$.
We have now completed the description of MNN. For the reader’s convenience, we provide a pseudocode of our algorithm in Appendix G.
4 Prediction Error Guarantee
We analyze the performance of MNN in terms of its prediction error at a user-item level. To that end, our analysis relies on three key ingredients: First, we show in that the our joint clustering &\&& denoising procedure yields good estimates for the outcomes of the central user-item pairs whenever they are revealed (see Proposition 4.1); Second, we characterize the pattern of the revealed central user-item pairs (see Proposition 4.2); Third, we show that the imputed central user-item pairs are accurate (4.3). In conclusion, we provide a finite-sample guarantee on the prediction error of MNN in Theorem 4.1, and discuss the obtained rates in comparison with the optimal parametric rates for supervised learning parametric.
4.1 Analysis
Clustering &\&& Denoising.
Let us denote the set of revealed user-item clusters as follows
$\Omega^{\star}=\{(i,j):i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star},\quad|\Omega_{ij}|\geq N_{n}\},$ (12)
In Proposition 4.1, we make the choice of $N_{n}$ precise and provide a finite-sample guarantee on the error rate between the revealed entries of $\overline{\mathbf{H}}$ and $\mathbf{H}$.
Proposition 4.1.
Let $\delta\in(0,1)$. Suppose we choose $N_{n}\gtrsim n^{2}\rho_{n}\varepsilon_{n}^{2d-2}$ then the event: $\forall(i,j)\in\Omega^{\star},$
$|\bar{h}_{ij}-h_{ij}|\lesssim LB\varepsilon_{n}+\sqrt{\frac{\sigma^{2}}{n^{2}\rho_{n}}\!\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\!\log\left(\frac{e\,n}{\delta}\right)},$ (13)
holds with probability at least $1-\delta$, provided $n\gtrsim d\log\left(e\,n/\delta\right)$ and $\varepsilon_{n}$ satisfies
$\varepsilon_{n}\gtrsim\frac{1}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)}.$ (14)
The precise statement of Proposition 4.1 is given in Lemma C.1, and its proof in Appendix C.
Observations Pattern of $\overline{\mathbf{H}}$.
In order to ensure correctness of the subroutine $\texttt{Impute}(\overline{\mathbf{H}},i,j)$, the mappings $\theta(\cdot)$ and $\phi(\cdot)$ must be “non-degenerate” in some appropriate sense so that the search for $\mathcal{I}$ and $\mathcal{J}$ is successful, and that the smallest singular value of $\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}$ is bounded away from 00 independently from $n$. This will in turn ensure that observation bias may still be exploited to impute the missing entries. Below, we precisely state what we mean by “non-degenerate”.
Definition 4.1 (Non-degeneracy).
We say that a mapping $\theta(\cdot)$ from $\mathcal{S}^{d-1}$ to $\mathbb{R}^{d}$ is non-degenerate with parameters $m,\kappa>0$ if it satisfies the following two properties: (a) for all $\mathbf{x}\in\mathcal{S}^{d-1}$, $\|\theta(\mathbf{x})\|_{2}\geq m\|\mathbf{x}\|_{2}$; (b) for all $\mathbf{x},\mathbf{y}\in\mathcal{S}^{d-1}$, $\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{y})\|_{2}\geq\kappa\|\mathbf{x}-\mathbf{y}\|_{2}$, where we define $\tilde{\theta}({\mathbf{x}})=\lVert\theta(\mathbf{x})\rVert^{-1}\theta(\mathbf{x})$.
Now, we are ready to present Proposition 4.2, in which we describe the pattern of revealed entries in $\overline{\mathbf{H}}$ provided that $\theta(\cdot)$ and $\phi(\cdot)$ are “non-degenerate”. This in turn ensures that $\texttt{Impute}(\overline{\mathbf{H}},i,j)$ will successfully terminate with high probability.
Proposition 4.2.
Assume that $\theta(\cdot)$ and ϕ()italic-$\phi(\cdot)$ are non-degenerate with parameters $m,\kappa>0$, and that $N_{n}$ is chosen as in Proposition 4.1. Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ be such that $\bar{h}_{i,j}=\star$, (i.e., $(i,j)$ is a non-revealed user-item cluster in $\overline{\mathbf{H}}$). Then, for all $\delta\in(0,1)$, the event: there exists $\mathcal{I}\subset\mathcal{U}^{\star},\mathcal{J}\subseteq\mathcal{V}^{\star}$ with $|\mathcal{I}|=|\mathcal{J}|=d$, such that
[…]
holds with probability at least $1-\delta$, provided that $\varepsilon_{n}$ satisfies
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}$ $\gtrsim\frac{1}{n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right),$
$\varepsilon_{n}$ $\lesssim\left(\frac{\kappa^{4}(m^{2}\wedge 1)}{L(B\vee 1)}\wedge 1\right).$ (16)
The proof of Proposition 4.3 is given in Appendix D. The precise statement of the proposition is given in D.1. It is worth noting that the “non-degeneracy” property allows us to rule out pathological cases for which it is impossible to impute the missing entries from the observed one, at least using our approach. The need for such property is apparent in the proof, and we refer the reader to Appendix E to see how this property comes into our analysis.
Matrix completion.
Next, we present a finite-sample error bound on the imputation error of all entries including the missing ones in $\overline{\mathbf{H}}$.
Proposition 4.3.
Let $\delta\in(0,1)$. Then, under the same assumption as in Proposition 4.2, the following event: $\forall(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$,
$|\hat{h}_{ij}-h_{ij}|\lesssim LB\varepsilon_{n}+\sqrt{\!\frac{\sigma^{2}}{n^{2}\rho_{n}}\!\!\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\!\!\!\!\log\left(\frac{e\,n}{\delta}\right)}$ (17)
holds with probability at least $1-\delta$, where $c=\textup{poly}(d,B,L,m^{-1},\kappa^{-1})$ provided that $n\gtrsim d\log(e\,n/\delta)$, and $\varepsilon_{n}$ satisfies (89), and
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}$ $\!\!\gtrsim\frac{d}{m^{2}\kappa^{4}n}\!\left(\!\!\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\!\!\right),$
$\varepsilon_{n}$ $\lesssim\left(\frac{\kappa^{4}(m^{2}\wedge 1)}{dL(B\vee 1)}\wedge 1\right).$ (18)
The precise statement of Proposition 4.3 is given in Lemma D.2 and its proof in presented in Appendix D.
Prediction error.
Finally, we are ready to present the main performance guarantee satisfied by MNN. To simplify the exposition of our result, we will consider that $\rho_{n}=n^{-\beta}$ for some $\beta>0$.
Theorem 4.1.
Suppose we are under the same assumptions as in Proposition 4.2. Let $\delta\in(0,1)$. Suppose we choose $\varepsilon_{n}=2n^{-\frac{2-\beta}{d-1}}$. Then, the event: $\forall(i,j)\in[n]\times[n]$
$|\hat{x}_{ij}-x_{ij}|\leq K\,n^{-\frac{2-\beta}{2d}}\sqrt{\log\left(\frac{e\,n}{\delta}\right)},$ (19)
holds with probability at least $1-\delta$, provided that $d>9/4$, $\beta<1/4$, and
$n\geq\left(\frac{CLBd^{3}}{m^{2}\kappa^{4}}\log\left(\frac{e\,n}{\delta}\right)\right)^{2d},$ (20)
where the constant $K=\textup{poly}(d,L,B,\sigma,m^{-1},\kappa^{-1})$, and $C>0$ is a universal constant.
The proof of Theorem 4.1 builds upon the previously stated propositions and its proof is deferred to Appendix F.
Observe that condition (20) in Theorem 4.1 can be verified [^1] This can be verified by using the principle branch of the Lambert function $W_{0}$, that satisfies $W_{0}(x)=y$ with $x=y^{y}$ for $x\geq 0$. Additionally, we have $W_{0}(x)\geq\log(x)-\log\log(x)+\frac{\log\log(x)}{\log(x)}$, for all $x\geq e$. (see Hoorfar and Hassani (2008)). if $d=O(\log(n)-\log(\log(n)))$ where $O(\cdot)$ may hide polynomial dependencies on $L,B,m^{-1}$ and $\kappa^{-1}$.
4.2 Discussion
Parametric rate in supervised learning.
Suppose we knew exactly the latent features $(\mathbf{u}_{i})_{i\in[n]}$, $(\mathbf{v}_{j})_{j\in[n]}$, then predicting outcomes of interest from noisy observations amounts to solving a supervised learning task where we recall that $f:\mathbb{R}^{2d}\mapsto\mathbb{R}$ is a bounded Lipschitz function that is a priori unknown (we recall from section 2 that we adopt a model of the form $y_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})+\xi_{ij}$). In this classical supervised learning setting, we know, for example from Theorem 1 in Stone (1982), that the optimal achievable parametric prediction error rate is
$\max_{i,j\in[n]}|\hat{x}_{ij}-x_{ij}|=\widetilde{O}\left(N^{-\frac{1}{2+2d}}\right)$ (21)
where $N$ would correspond to the number of observations and $\widetilde{O}(\cdot)$ may hide logarithmic dependencies on $N$. Let us further add that the $k$-nearest neighbor is an example of supervised learning method that achieve such rate Chen and Shah (2018).
Comparison with the parametric rate.
The latent vectors $(\mathbf{u}_{i})_{i\in[n]}$, $(\mathbf{v}_{j})_{j\in[n]}$ in our proposed model are not observed as they represent confounding factors. Despite this additional challenge, the proposed algorithm MNN still attains a competitive rate in comparison with the parametric one for the supervised learning setting described above. Indeed, we provide the following result:
Lemma 4.1.
Let $N$ denote the expected number of observed noisy entries of $\mathbf{X}$. Suppose we are under the same assumptions as in Theorem 4.1, and with similar choices of $N_{n}$ and $\varepsilon_{n}$. Then for all $\delta\in(0,1)$, the event:
$\max_{i,j\in[n]}|\hat{x}_{ij}-x_{ij}|=\widetilde{O}\left(N^{-\frac{1}{2d}}\sqrt{\log\left(\frac{e}{\delta}\right)}\right)$ (22)
holds with probability at least $1-\delta$, provided similar conditions as in Theorem 4.1 hold, and where $\widetilde{O}(\cdot)$ hides logarithmic dependencies on $\log(N)$ and polynomial dependencies in $d,L,B,\sigma,m^{-1},$ and $\kappa^{-1}$.
To spare the reader unnecessary details, we omit the proof of Lemma 4.1 as the result follows immediately form Theorem 4.1. To see that, note that under the proposed model, according to (2), the expected number of observation $N\approx n^{2-\beta}$. Rewriting this as $n=N^{-1/(2-\beta)}$ and plug-in it in the error rate expression (19) in Theorem 4.1 yields the desired result.
In conclusion, estimating $x_{ij}$ using MNN achieves a prediction error rate that is competitive with the parametric rate for sufficiently slowly decaying $\rho_{n}$, despite the following unconventional challenges: (i) the algorithm does not have direct access to the latent features $(\mathbf{u}_{i})_{i\in[n]}$ and $(\mathbf{v}_{i})_{i\in[n]}$, but instead has to estimate them; (ii) samples are drawn in a biased manner as opposed to the uniform sampling of datapoints over the domain, which is typically assumed.
5 Experiments
5.1 Synthetic Data
Data.
We performed several experiments with different numbers of datapoints $n$ and sparsity values $\rho_{n}$, all of which had data generated using the following steps. First, we generated rank $d=5$ latent factors $\mathbf{U}\in[-1,1]^{n\times d}$, $\mathbf{V}\in[-1,1]^{n\times d}$ by sampling $d$ entries i.i.d from a standard normal distribution for each of the $n$ users and items and then normalizing each row to have unit norm. We then generated the matrix $\mathbf{P}$ of probabilities as described in our observation model by setting $\mathbf{P}=\frac{\rho_{n}}{2}(\mathbf{U}\mathbf{V}^{\top}+\mathbf{1}\mathbf{1}^{\top})$. Next, we sample $\mathbf{A}\overset{\operatorname{iid}}{\sim}\mathsf{Ber}(\mathbf{P})$. We defined our outcome matrix $\mathbf{X}=\theta(\mathbf{U})\phi(\mathbf{V})^{\top}$ where
$\theta(\mathbf{U})=e^{\sqrt{d}\mathbf{U}}$ (23)
and
$\phi(\mathbf{V})=e^{\sqrt{d}\mathbf{V}}.$ (24)
We then generated the subgaussian noise terms $\xi_{ij}$ by sampling $n^{2}$ entries i.i.d from a standard normal distribution. Finally, the partially-observed noisy outcome matrix $\mathbf{Y}$ was generated by setting $y_{ij}=x_{ij}+\xi_{ij}$ if $a_{ij}=1$ and $y_{ij}$ to be a null value otherwise.
Algorithms compared.
We compare the performance of the proposed algorithm against a standard matrix completion algorithm USVT. In the experiments, we utilize a practical implementation of the proposed MNN algorithm where (i) $\mathbf{A}$ is used in place of $\tilde{\mathbf{A}}$ in the distance estimation step and (ii) the matrix completion from the clustered outcome matrix is done using SVD with known rank. Using such a practical implementation of the algorithm acts a proof of concept – letting us determine the added value of the extra information gained from the two stage approach. We compare against a modified version of USVT where the rank of the matrix is known, allowing for more comparable performance against MNN which utilizes the matrix rank. We note that MNN is designed to handle biased data while the modified USVT algorithm is not.
Metrics.
We use $R^{2}$ score, MSE, MAE, and max error to assess the similarity of the estimated full outcome matrix with the true outcome matrix. We directly compare the distribution of estimated entries to the distribution of true outcomes to assess the bias of the algorithms.
Experiment setup.
We performed two different experiments with the described algorithms, using 16-fold cross-validation to tune the hyperparameters. In the first experiment, we tried 3 values of $n\in\{100,562,3162\}$. We set $\rho_{n}=n^{-1/4}$. We then ran both MNN and modified USVT with these values of $n$ and compared the resulting estimates. In the second experiment, we set $n=1000$ and tested 3 values of $\beta\in\{0,1/6,1/4\}$. We set $\rho_{n}=n^{-\beta}$. We then again ran both MNN and modified USVT and compared the bias of the resulting estimates. We concluded by running 10 experimental repeats of MNN and modified USVT using data generated with different random seeds for each of these $\rho_{n}=n^{-\beta}$ values. These final runs were used to calculate the $R^{2}$, MSE, MAE, and max error as a function of $n$ for each $\beta$.
Results.
Before comparing the performance of MNN and modified USVT on the synthetic dataset, we examine the bias of the estimates for the full outcome matrix in both cases (experiments 1 and 2). As can be seen in Fig. 1,the distribution of estimates generated by MNN better approximates the true distribution of outcomes as the number of datapoints increases. Moreover, as $n$ increases, the shape of the distribution indicates that the results are less biased. However, the estimates generated by modified USVT do not display this same trend and do not appear to become less biased as the number of datapoints increases. This difference in bias can also be seen in Fig. 2 as $n$ is held constant at 1000. As $\rho_{n}$ decreases from 1 to $n^{-1/4}$, the bias of MNN barely changes but the bias of modified USVT increases significantly.
Figure 1: Distributions for synthetic data generated with $n=100,562$, and $3162$ users and items (top to bottom).
Figure 2: Predictions compared to true values for MNN (left) and modified USVT (right) with $\beta=0,\beta=1/6$, and $\beta=1/4$ (top to bottom).
Now, we consider $R^{2}$ score, MSE, MAE, and max error as metrics to compare the estimates made by MNN and modified USVT against the true outcomes. The results of this experiment can be seen in Table 1. We can see that across all these metrics, MNN outperforms modified USVT. When comparing the $R^{2}$ scores, we see that the difference in the performance of the two algorithms remains relatively constant for different $\rho_{n}$ values. The results of these $R^{2}$-, MSE-, MAE-, and max error-based metrics all line up with the observations made about the relative bias of the estimates in the two algorithms, indicating that the latent factor clustering approach utilized by MNN is indeed de-biasing the data and leads to better estimates.
Algorithm $\beta=0$ $\beta=1/6$ $\beta=1/4$
$R^{2}$ MNN .$\mathbf{.831\pm.003}$$\mathbf{.573\pm.011}$$\mathbf{.394\pm.032}$
USVT .$393\pm.003$$200\pm.017$$-.015\pm.014- .$
MSE MNN $\mathbf{14.1\pm.291}$ $\mathbf{35.6\pm.900}$ $\mathbf{50.5\pm 2.67}$
USVT $50.6\pm.250$ $66.7\pm 1.42$ $84.7\pm 1.19$
MAE MNN $\mathbf{2.62\pm.028}$ $\mathbf{4.34\pm.066}$ $\mathbf{5.37\pm.168}$
USVT $4.61\pm.008$ $5.05\pm.043$ $5.57\pm.037$
Max Error MNN $\mathbf{36.0\pm 2.24}$ $\mathbf{56.3\pm 3.56}$ $\mathbf{65.6\pm 2.65}$
USVT $42.0\pm.616$ $59.3\pm 2.20$ $66.5\pm 1.44$
Table 1: $R^{2}$, MSE, MAE, and max error for matrix completion methods on synthetic datasets.
5.2 Real-World Data
Data.
We used a dataset obtained from Glance [^2] Glance (www.glance.com) is a smart lock screen that aims to personalize the screen. detailing the interaction of users with different pieces of content, considering a partially observed matrix where rows represent users, columns represent the different pieces digital content, and outcomes are the natural logarithm of the duration (in seconds) that the user interacted with the content. The dataset we considered had 1014696 observed measurements for 1305 users and 1471 content items. The sparsity of the dataset was thus $53\%$ and consistent with the observation model described in our setup.
Algorithms compared.
We compare the same practical implementation of MNN and a modified version of USVT described in the synthetic data experiments.
Metrics.
As in the synthetic data experiments, we used $R^{2}$ score, MSE, MAE, and max error to assess the similarity of our estimates with the true outcomes for the test sets. We also directly compare the distributions of estimated entries to true outcomes in our test set.
Experiment setup.
We split the data into $90/10$ train/test sets 10 times. We determined the best estimate of the rank of the true outcome matrix and the rank of the observation pattern using 9-fold cross-validation with MNN. We used 16-fold cross-validation to separately tune the other hyperparameters. Next, we ran both MNN and modified USVT on the training set in each experimental run. Using the tuned hyperparameters, we ran both algorithms for 10 experimental runs using different train/test splits to calculate performance metrics.
Results.
As we can see from Fig. 4, the estimates from modified USVT are extremely biased when the algorithm is used with MNAR data. The estimates from MNN, however, appear to be minimally biased. Moreover, from Fig. 3, we can see that the estimates made by modified USVT are very sensitive to outliers in the data, while the estimates from MNN are not. We also observe from Table 2 that the $R^{2}$, MSE, MAE, and max error of MNN are far better than that of modified USVT, implying that MNN works significantly better on MNAR data in practice.
Figure 3: Distributions of test set estimates from MNN and modified USVT for real-world data.
Figure 4: Predictions vs true outcomes for real-world data using MNN (left) and modified USVT (right).
Algorithm $R^{2}$ MSE MAE Max Error
MNN .$\mathbf{.039\pm.038}$ $\mathbf{1.15\pm.054}$$\mathbf{.843\pm.021}$ $\mathbf{6.49\pm.405}$
USVT $-26.0\pm.435$ $32.4\pm.392$ $4.95\pm.032$ $11.6\pm.053$
Table 2: $R^{2}$, MSE, MAE, and max error for matrix completion methods on real data.
Appendix A Distance Estimation
To guarantee good clustering performance, we must first show that the estimated distances converge to the true distances between pairs of users/items.
Lemma A.1 (Distance estimation).
For all $\delta\in(0,1)$, we have
$\mathbb{P}\left(\max_{i\neq j}\left|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}\right|\leq\frac{c}{\rho^{2}_{n}}\sqrt{\frac{d^{3}}{n}\log\left\lparen\frac{e\,n}{\delta}\right\rparen}\right)\geq 1-\delta,$ (25)
provided that $n\geq Cd\log(en/\delta)$, for some universal constants $c,C>0$.
Proof of Lemma A.1.
Let $i,j\in[n]$, and let us recall that
$\hat{d}^{\mathbf{u}}_{ij}=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widehat{\widetilde{\mathbf{P}}}_{j,:}\right\rVert_{2}\qquad\text{and}\qquad d^{\mathbf{u}}_{ij}=\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}.$ (26)
For ease of notations, let us further introduce the distance corresponding to the case where we know exactly the true observation probability matrix $\mathbf{P}$,
$\tilde{d}^{\mathbf{u}}_{ij}=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-{\widetilde{\mathbf{P}}}_{j,:}\right\rVert_{2}.$ (27)
Note that we can decompose the estimation error as follows using the triangular inequality
$\left|\hat{d}^{\mathbf{u}}_{ij}-d^{\mathbf{u}}_{ij}\right|\leq\left|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}\right|+\left|\tilde{d}^{\mathbf{u}}_{ij}-d^{\mathbf{u}}_{ij}\right|.$ (28)
We will next proceed by bounding each term separately with high probability.
(1) Bounding $|\tilde{d}^{\mathbf{u}}_{ij}-d^{\mathbf{u}}_{ij}|$. We start by noting that
$\tilde{d}_{ij}^{\mathbf{u}}=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}$ $=\sqrt{\frac{d}{n}}~{}\left\lVert\mathbf{V}(\mathbf{u}_{i}-\mathbf{u}_{j})\right\rVert_{2}.$ (29)
Thus,
[…]
which entails that
$\left|\tilde{d}_{ij}^{\mathbf{u}}-d^{\mathbf{u}}_{ij}\right|=\left|\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}-\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}\right|\leq\max_{k\in[d]}\left|\sqrt{\frac{d}{n}}~{}\sigma_{k}\left(\mathbf{V}\right)-1\right|\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}$ (31)
where $\sigma_{\min}(\mathbf{X})$ (resp. $\sigma_{\max}(\mathbf{X})$) refers to the smallest (resp. largest) singular value of $\mathbf{X}$. Since $\sqrt{d}\cdot\mathbf{V}$ is a random matrix with i.i.d. isotropic rows, using Lemma H.5 and picking $t=\sqrt{\log(2/\delta)/c}$, we obtain that
[…]
holds with probability at least $1-\delta$, for some universal constant $C>0$. Thus, recalling the simple fact $\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}\leq 2$, we conclude that
$\mathbb{P}\left(\left|\tilde{d}_{ij}^{\mathbf{u}}-d^{\mathbf{u}}_{ij}\right|\leq C\sqrt{\frac{d}{n}\log\left\lparen\frac{2}{\delta}\right\rparen}\right)\geq 1-\delta.$ (33)
(2) Bounding $|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}|$. Using first the reverse triangular inequality, then the triangular inequality, we can further decompose the error term $|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}|$ as follows
$\left|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}\right|$ $\leq\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}\left\|(\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widehat{\widetilde{\mathbf{P}}}_{j,:})-(\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:})\right\|_{2}$ (34)
$\leq\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}\left(\left\|(\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\right\|+\left\|\widehat{\widetilde{\mathbf{P}}}_{j,:}-\widetilde{\mathbf{P}}_{j,:}\right\|_{2}\right)$ (35)
Therefore controlling the term $|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}|$ amounts to showing that $\|\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\|_{2}$ is small for all $i\in[n]$. To that end, we use Lemma H.2 to write
$\left\|\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\right\|_{2}^{2}\leq 2\cdot\left(\frac{\|\mathbf{E}\|_{\textup{op}}^{2}}{\sigma_{d}^{2}(\widetilde{\mathbf{P}})}\left(\|\mathbf{E}_{j,:}\|_{2}^{2}+\|\widetilde{\mathbf{P}}_{j,:}\|_{2}^{2}\right)+\|\mathbf{R}_{d}\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\|_{2}^{2}\right),$ (36)
where we denote $\mathbf{E}=\widetilde{\mathbf{A}}-\widetilde{\mathbf{P}}=\mathbf{A}-\mathbf{P}$, and $\mathbf{Q}_{d}\mathbf{\Sigma}_{d}\mathbf{R}_{d}^{\top}$ the reduced SVD of $\widetilde{\mathbf{P}}$. Now, we will produce several bounds on the norms of the relevant matrices appearing in (36) separately.
(2.1) Upper bounding $\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}$. The entries of the noise matrix $\mathbf{E}$ are independent centered Bernoulli random variables, thus sub-gaussian. Using Lemma H.3 and choosing $t=\sqrt{\log(2/\delta)/c}$ again, we obtain that the following
$\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}$ $\leq 2C\sqrt{n}+\sqrt{\frac{\log(2/\delta)}{c}}\leq C^{\prime}\sqrt{n\log(2/\delta)}.$ (37)
holds with probability at least $1-\delta$, for some universal constants $c,C,C^{\prime}>0$.
(2.2) Lower bounding $\sigma_{d}(\widetilde{\mathbf{P}})$. We recall $\widetilde{\mathbf{P}}=(\rho_{n}/2)\mathbf{U}\mathbf{V}^{\top}$. By (32), the following
$\sigma_{d}(\mathbf{V})$ $\geq\sqrt{\frac{n}{d}}-C^{\prime}\sqrt{\log{\left\lparen\frac{2}{\delta}\right\rparen}}$ (38)
$\geq C^{\prime\prime}\sqrt{\frac{n}{d}}.$ (39)
holds with probability $1-\delta$, provided that $n\geq(2C^{\prime})^{2}d\log(2/\delta)$ for some universal constants $C^{\prime},C^{\prime\prime}>0$. By the same argument, we obtain a similar high probability lower bound on $\sigma_{d}(\mathbf{U})$. Then, by a simple union bound, we obtain that
$\sigma_{d}(\widetilde{\mathbf{P}})$ $\geq\frac{\rho_{n}}{2}\cdot\sigma_{d}(\mathbf{U})\cdot\sigma_{d}(\mathbf{V})$ (40)
$\geq\frac{C\rho_{n}n}{d},$ (41)
holds with probability at least $1-\delta$, provided $n\geq cd\log(2/\delta)$, for some universal constants $c,C>0$.
(2.3) Upper bounding $\left\lVert\mathbf{E}_{j,:}\right\rVert_{2}$. Note that $\|\mathbf{E}_{j,:}\|_{2}\leq\|\mathbf{E}\|_{{\textup{op}}}$. Thus, we have
$\left\lVert\mathbf{E}_{j,:}\right\rVert_{2}\leq C^{\prime}\sqrt{n\log(2/\delta)}$ (42)
holds with probability at least $1-\delta$, as shown in (37).
(2.4) Upper bounding $\left\lVert\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}$. Recall that $\widetilde{\mathbf{P}}_{j,:}=(\rho_{n}/2)\mathbf{V}\mathbf{u}_{j}$. Using (32), which bounds the largest singular value of $\mathbf{V}$, we have that
$\left\lVert\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}$ $=\frac{\rho_{n}}{2}\left\lVert\mathbf{V}\mathbf{u}_{j}\right\rVert_{2}$ (43)
$\leq\frac{\rho_{n}}{2}\left\lVert\mathbf{V}\right\rVert_{\operatorname{op}}\left\lVert\mathbf{u}_{j}\right\rVert_{2}$ (44)
[…]
$\leq\frac{C^{\prime\prime}\rho_{n}}{2}\sqrt{\frac{n}{d}}.$ (46)
holds with probability at least $1-\delta$, provided $n\geq c^{\prime}d\log(2/\delta)$.
(2.5) Upper bounding $\left\lVert\mathbf{R}_{d}\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\right\rVert_{2}$. First, since $\mathbf{R}_{d}$ contains orthonormal columns,
$\left\lVert\mathbf{R}_{d}\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\right\rVert_{2}$ $\leq\left\lVert\mathbf{R}_{d}\right\rVert_{\operatorname{op}}\left\lVert\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\right\rVert_{2}$ (47)
$\leq\left\lVert\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\right\rVert_{2}.$ (48)
We will bound this quantity using Lemma H.7. Towards showing that $\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}$ is a sub-gaussian random vector, consider any fixed $\mathbf{v}\in\mathcal{S}^{d-1}$. Then, denoting $\mathbf{R}_{d}=[r_{ki}]_{k\in[n],i\in[d]}$,
$\mathbf{v}^{\top}\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}=\sum_{i=1}^{d}\sum_{k=1}^{n}v_{i}r_{ki}\eta_{jk}$ (49)
which is a linear combination of independent sub-gaussian random variables $\eta_{jk}\sim\,\mathsf{subG}\,(1/4)$, and thus sub-gaussian. Moreover, its variance proxy is
$\frac{1}{4}\sum_{i=1}^{d}\sum_{k=1}^{n}v_{i}^{2}r_{ki}^{2}$ $=\frac{1}{4}\sum_{i=1}^{d}\left\lparen v_{i}^{2}\sum_{k=1}^{n}r_{ki}^{2}\right\rparen$ ) (50)
$=\frac{1}{4}\sum_{i=1}^{d}v_{i}^{2}$ (51)
$=\frac{1}{4}.$ (52)
Therefore, $\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}$ is a $(1/4)$-sub-gaussian random vector. Using Lemma H.7 with $t=\sqrt{\frac{d}{2}\log\left\lparen\frac{2n}{\delta}\right\rparen}$ and union bounding over all rows of $\mathbf{E}$ yields
$).\max_{j\in[n]}~{}\left\lVert\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\right\rVert_{2}\leq\sqrt{\frac{d}{2}\log\left\lparen\frac{2n}{\delta}\right\rparen}.$ (53)
with probability at least $1-\delta$.
(2.6) Putting everything together. Starting from (36), and using the bounds obtained in (37), (40), (42), (43), and (53), we have by substitution that
[…]
$\leq C\cdot\left\lparen\frac{C^{\prime}d^{2}}{\rho_{n}^{2}}+\frac{d}{2}\log\left\lparen\frac{2n}{\delta}\right\rparen\right\rparen$ ) ) (56)
$\leq\frac{Cd^{2}}{\rho_{n}^{2}}\log\left\lparen\frac{2n}{\delta}\right\rparen.$ ) . (57)
Scaling appropriately according to (26), and using the union bound, then yields
$\mathbb{P}\left(\left|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}\right|\leq\frac{C}{\rho^{2}_{n}}\sqrt{\frac{d^{3}}{n}\log\left\lparen\frac{2n}{\delta}\right\rparen}\right)\geq 1-10\delta$ (58)
provided $n\geq cd\log(2/\delta)$ for some universal constants $c,C>0$.
(3) Conclusion. Finally, combining (33) and (58) with further simplifications gives
$\mathbb{P}\left(\max_{i\neq j}\left|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}\right|\leq\frac{C}{\rho^{2}_{n}}\sqrt{\frac{d^{3}}{n}\log\left\lparen\frac{en}{\delta}\right\rparen}\right)\geq 1-\delta,$ (59)
provided $n\geq cd\log(en/\delta)$, for some universal constants $c,C>0$. ∎
Appendix B Clustering
Coverage. A key phenomenon that underpins our results is dense coverage of the latent feature space by users and items. Towards guaranteeing this, let $\Sigma_{n}\subset\mathcal{S}^{d-1}$ be a minimal $\varepsilon_{n}$-net of $\mathcal{S}^{d-1}$, with the properties that $\left\lVert\mathbf{z}-\mathbf{z}^{\prime}\right\rVert_{2}>\varepsilon_{n}$ for distinct $\mathbf{z},\mathbf{z}^{\prime}\in\Sigma_{n}$ and for any $\mathbf{x}\in\mathcal{S}^{d-1}$ there exists $\mathbf{z}\in\Sigma_{n}$ such that $\left\lVert\mathbf{x}-\mathbf{z}\right\rVert_{2}\leq\varepsilon_{n}$. It is widely known that $|\Sigma_{n}|\leq(1+2/\varepsilon_{n})^{d}\leq(3/\varepsilon_{n})^{d}$ where the latter inequality assumes $\varepsilon_{n}\leq 1$, which we enforce.
Lemma B.1 (Coverage).
Denote the $\varepsilon_{n}$-ball centered at $\mathbf{z}$ as $\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})$, and let $(\mathbf{u}_{i})_{i\in[n]}$ be independent and uniformly random points on $\mathcal{S}^{d-1}$. Then, for all $\delta\in(0,1)$,
$\mathbb{P}\left(\min_{\mathbf{z}\in\Sigma_{n}}\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\geq\frac{n}{8}\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right)\geq 1-\delta,$ (60)
provided that $\varepsilon_{n}$ satisfies
$n\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\geq C\log\left(\frac{en}{\delta}\right),$ (61)
where $C$ is a positive universal constant.
Proof of Lemma B.1.
First, let us fix $\mathbf{z}\in\Sigma_{n}$. Note that by assumption $\mathbf{u}_{1},\dots,\mathbf{u}_{n}$ are i.i.d. on the unit sphere, so and an immediate application of multiplicative Chernoff bound with ratio 1/2 gives
$\mathbb{P}\left(\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{2}\,\mathbb{P}(\mathbf{u}_{1}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\right)\leq\exp\left(-\frac{n}{8}\,\mathbb{P}(\mathbf{u}_{1}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\right)$ (62)
Now, by using Lemma H.1 we have
$\mathbb{P}(\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\geq\frac{1}{4}\sin^{d-1}\Phi$ (63)
where, using the geometry of the problem and recalling that $\varepsilon_{n}\leq 1$,
$\sin\Phi=\varepsilon_{n}\sqrt{1-\varepsilon_{n}^{2}/4}\geq\varepsilon_{n}/2.$ (64)
Thus we can write
$\mathbb{P}(\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\geq\frac{1}{4}\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}.$ (65)
This entails that
$\mathbb{P}\left(\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right)\leq\exp\left(-\frac{n}{32}\,\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right).$ ) . (66)
Finally, by using an union bound over the $\varepsilon_{n}$-net $\Sigma_{n}$, we conclude that
$\mathbb{P}\left(\min_{\mathbf{z}\in\Sigma_{n}}\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right)$ $\leq\left(\frac{3}{\varepsilon_{n}}\right)^{d}\exp\left(-\frac{n}{32}\,\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right)$ ) (67)
$\leq\delta,$ (68)
where the last inequality is true whenever the condition $n\left(\varepsilon_{n}/2\right)^{d-1}\geq 32\left(\log\left((1/\delta)\right)+d\log\left(3/\varepsilon_{n}\right)\right),$ holds, which we may replace by the more conservative condition
$n\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\geq 32\log\left(\frac{en}{64\delta}\right).$ (69)
Observations. Given that the latent feature space is densely covered by both users and items, we now argue that all pairs of user and item clusters – as defined by $\Sigma_{n}$, rather than the central users and items – that are reasonably close will generate a guaranteed number of observations.
Lemma B.2 (Observations).
Consider $\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}$ for which $\left\lVert\mathbf{z}_{i}-\mathbf{z}_{j}\right\rVert_{2}\leq\sqrt{2}+c\varepsilon_{n}$ for some absolute constant $c>0$, where we assume that $\varepsilon_{n}<(\sqrt{3}-\sqrt{2})/(2+c)$, and let the set of observations between these user and item clusters be denoted by
$\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})=\{y_{kl}:a_{kl}=1,\mathbf{u}_{k}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i}),\mathbf{v}_{l}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j})\}.$ (70)
Then, for all $\delta\in(0,1)$,
$\mathbb{P}\left(\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})\right\rvert\geq cn^{2}\rho_{n}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right)\geq 1-\delta,$ (71)
as long as $\varepsilon_{n}$ satisfies
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\geq\frac{C}{n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right)$ (72)
for some universal constants $c,C>0$.
Proof of Lemma B.2.
Let us define the events
$\mathcal{E}_{1}$ =$=\left\{\sum_{k=1}^{n}\mathds{1}\{\mathbf{u}_{k}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i})\}\geq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right\},$ (73)
$\mathcal{E}_{2}$ =$=\left\{\sum_{\ell=1}^{n}\mathds{1}\{\mathbf{v}_{\ell}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j})\}\geq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right\},$ (74)
and note that according to Lemma B.1 and union bounding over the two events we have
$\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})\geq 1-2\delta$
provided that $1\geq\varepsilon_{n}\geq 2\left(\frac{C}{n}\log\left(\frac{en}{\delta}\right)\right)^{\frac{1}{d-1}}$. Next, observe that for any $\mathbf{u}_{k}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i}),\mathbf{v}_{l}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j})$ where $\left\lVert\mathbf{z}_{i}-\mathbf{z}_{j}\right\rVert_{2}\leq\sqrt{2}-2\varepsilon_{n}$, we have
$\left\lVert\mathbf{u}_{k}-\mathbf{v}_{l}\right\rVert_{2}\leq\left\lVert\mathbf{z}_{i}-\mathbf{z}_{j}\right\rVert_{2}+\left\lVert\mathbf{u}_{k}-\mathbf{z}_{i}\right\rVert_{2}+\left\lVert\mathbf{v}_{l}-\mathbf{z}_{j}\right\rVert_{2}\leq\sqrt{2}+(2+c)\varepsilon_{n}\leq\sqrt{3}.$ (75)
Thus,
$p_{kl}$ $=\frac{\rho_{n}}{2}(\mathbf{u}_{k}^{\top}\mathbf{v}_{l}+1)$ (76)
$=\frac{\rho_{n}}{2}\left\lparen 2-\frac{1}{2}\left\lVert\mathbf{u}_{k}-\mathbf{v}_{l}\right\rVert_{2}^{2}\right\rparen$ ) (77)
$\geq\rho_{n}/4.$ (78)
Now, conditioned on the event $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, we know that there are at least $\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}$ users and items in the balls around $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ respectively, all pairs of which satisfy the above inequality. Defining
$\mathcal{E}_{3}=\left\{\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})\right\rvert\geq\frac{n^{2}\rho_{n}}{512}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right\},$ (81)
then using a multiplicative Chernoff bound and union bounding over all pairs of points in $\Sigma_{n}$, we get
[…]
where the last inequality holds as long as
$\rho_{n}n^{2}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\geq 2048\left(\log\left(\frac{1}{\delta}\right)+2d\log\left(\frac{3}{\varepsilon_{n}}\right)\right).$ (83)
Note that $n(\varepsilon_{n}/2)^{d-1}\geq C^{\prime}\log(en/\delta)$ ensures that $\log(n)\geq d\log(\varepsilon_{n}/3)$ for $C^{\prime}$ larger than $1$. Thus, the above condition can be enforced if,
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\geq\frac{C}{n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right)$ (84)
for some positive universal constants $C$ that are large enough. Finally, we conclude by noting that
$\mathbb{P}\left(\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})\right\rvert<\frac{\rho_{n}n^{2}}{512}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right)\leq\mathbb{P}\left(\mathcal{E}_{3}^{c}|\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)+\mathbb{P}(\mathcal{E}_{1}^{c}\cup\mathcal{E}_{2}^{c})\leq 3\delta.$ (87)
Appendix C Denoising
Here we argue that every revealed entry of the unfilled clustered outcome matrix $\bar{\mathbf{H}}$ is an accurate estimate of its corresponding entry in $\mathbf{H}$.
Lemma C.1 (Denoising).
Suppose we choose
$N_{n}=cn^{2}\rho_{n}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2},$
for some universal constants $c>0$ large enough. Then, for all $\delta\in(0,1)$, the following event: for every revealed entry $(i,j)$ of $\overline{\mathbf{H}}$
$\left\lvert\bar{h}_{ij}-h_{ij}\right\rvert\leq 14LB\varepsilon_{n}+\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\log\left(\frac{2n^{2}}{\delta}\right)}$ (88)
holds with probability at least $1-\delta$, provided that
$\varepsilon_{n}\geq\frac{C}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)},$ (89)
$n\geq Cd\log\left(\frac{e\,n}{\delta}\right),$ (90)
for some universal constant $C>0$.
Proof of Lemma C.1.
We can express the difference between $\bar{h}_{ij}$ and $h_{ij}$ for any $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ as
$\bar{h}_{ij}-h_{ij}$ $=\underbrace{\frac{1}{|\Omega_{ij}|}\sum_{(k,l)\in\Omega_{ij}}\left(f(\mathbf{u}_{k},\mathbf{v}_{l})-f(\mathbf{u}_{i},\mathbf{v}_{j})\right)}_{\Delta_{ij}}+\underbrace{\frac{1}{|\Omega_{ij}|}\sum_{(k,l)\in\Omega_{ij}}\xi_{kl}}_{\nu_{ij}}.$ (91)
(1) Bounding $\nu_{ij}$. Recall that an entry of $\overline{\mathbf{H}}$ is considered revealed if $\lvert\Omega_{ij}\rvert\geq N_{n}=c\,n^{2}\rho_{n}\left(\varepsilon_{n}/2\right)^{2d-2}$. Thus, $\nu_{ij}$ is the mean of at least $N_{n}$ independent subgaussian random variables of variance proxy $\sigma^{2}$, and by union bounding Lemma H.6 over all $n^{2}$ user-item pairs we have
$\mathbb{P}\left(\max_{(i,j):\bar{h}_{ij}\neq\star}\lvert\nu_{ij}\rvert\geq\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\log\left(\frac{2n^{2}}{\delta}\right)}\right)\leq\delta.$ (92)
(2) Bounding $\Delta_{ij}$. Moreover, by Lemma A.1, all the estimated user-user and item-item distances have error bounded by $\varepsilon_{n}$ if
$\delta_{\operatorname{dist}}(n)=\frac{C}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)}\leq\varepsilon_{n}.$
Since for any $k\in\mathcal{U}_{i}$ it must be true that $\hat{d}^{\mathbf{u}}_{ik}\leq 6\varepsilon_{n}$ by construction (otherwise user $k$ itself would have been a central user), we know the true distance is also small, i.e. $\left\lVert\mathbf{u}_{i}-\mathbf{u}_{k}\right\rVert_{2}\leq\hat{d}^{\mathbf{u}}_{ik}+\delta_{\operatorname{dist}}(n)\leq 7\varepsilon_{n}$. The same holds for $\left\lVert\mathbf{v}_{j}-\mathbf{v}_{l}\right\rVert_{2}$ for any $l\in\mathcal{V}_{j}$. Thus, by the Lipschitz property of $f$, every $(k,l)\in\Omega_{ij}$ satisfies
$\left\lvert f(\mathbf{u}_{k},\mathbf{v}_{l})-f(\mathbf{u}_{i},\mathbf{v}_{j})\right\rvert\leq 14LB\varepsilon_{n}.$ (93)
Combining the two bounds yields the desired bound. ∎
Appendix D Imputation
Pattern of observations and low-rank structure.
Once clustering &\&& denoising have been executed, the resulting matrix $\overline{\mathbf{H}}$ has still some missing entries, and its low-rank structure is a priori unclear. In Lemma D.1, we argue that $\overline{\mathbf{H}}$ has rank at least $d$, and describe the pattern its revealed entries. The theoretical tools we use towards proving Lemma D.1 exploit the connection between the geometry of the problem and the low-rank structure present in our model. These tools are presented in Appendix E.
Lemma D.1 (Patterns of observations).
Assume that $\theta(\cdot)$ and ϕ()italic-$\phi(\cdot)$ are non-degenerate with parameters $m,\kappa>0$. Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ be such that $\bar{h}_{i,j}=\star$, (i.e., $(i,j)$ is a non-revealed user-item cluster in $\overline{\mathbf{H}}$). Suppose we choose
$N_{n}=Cn^{2}\rho_{n}\varepsilon_{n}^{2d-2},$ (94)
where $C>0$ is an absolute constant sufficiently large. Then, there exists $\mathcal{I}\subset\mathcal{U}^{\star},\mathcal{J}\subseteq\mathcal{V}^{\star}$ with $|\mathcal{I}|=|\mathcal{J}|=d$, such that the event:
[…]
holds with probability at least $1-\delta$, provided that $\varepsilon_{n}$ satisfies
[…]
where $c,C,C^{\prime}>0$ are universal constants.
Proof of Lemma D.1.
Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ such that $\bar{h}_{ij}=\star$, and recall that $h_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})$. We start by invoking Proposition E.1, which ensures, following Algorithm 1, that there exists $\mathbf{x}_{1},\dots,\mathbf{x}_{d},$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$, such that:
  • (i)
    For all $k,\ell\in[d]$, $\|\mathbf{v}_{j}-\mathbf{x}_{k}\|_{2}\leq\sqrt{2}$, $\|\mathbf{u}_{i}-\mathbf{y}_{\ell}\|_{2}\leq\sqrt{2}$ and $\|\mathbf{x}_{k}-\mathbf{y}_{\ell}\|_{2}\leq\sqrt{2}$;
  • (ii)
    For all $i,j\in[d]$, $i\neq j$, we have $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}\wedge\|\mathbf{y}_{i}-\mathbf{y}_{j}\|_{2}\geq L^{-1}\kappa^{2}\sqrt{(1-\cos(\pi/12))\cos(\pi/12)}$;
  • (iii)
    And the following holds
    $\sigma_{d}\left(\begin{bmatrix}\theta(\mathbf{x}_{1})&\dots&\theta(\mathbf{x}_{d})\end{bmatrix}^{\top}\begin{bmatrix}\phi(\mathbf{y}_{1})&\ldots&\phi(\mathbf{y}_{d})\end{bmatrix}\right)\geq\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{2}$
Next, we argue that there exists =$\mathcal{I}=\{i_{1},\dots,i_{d}\}\subseteq\mathcal{U}^{\star}$ and $\mathcal{J}=\{j_{1},\dots,j_{d}\}\subseteq\mathcal{V}^{\star}$ such that
$\|\mathbf{x}_{k}-\mathbf{u}_{i_{k}}\|_{2}\leq 8\varepsilon_{n}\quad\text{and}\quad\|\mathbf{y}_{\ell}-\mathbf{v}_{j_{\ell}}\|_{2}\leq 8\varepsilon_{n}.$ (97)
Indeed, recall that $\Sigma_{n}$ is an $\epsilon_{n}$-net, thus there exists $\mathbf{z}_{1},\dots,\mathbf{z}_{d}\in\Sigma_{n}$ such that $\|\mathbf{z}_{k}-\mathbf{x}_{k}\|_{2}\leq\varepsilon_{n}$. Moreover thanks to property $(ii)$, satisfied by $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$, the vectors $\mathbf{z}_{1},\dots,\mathbf{z}_{d}$ are distinct, with every pair being $14\varepsilon_{n}$ apart if
$14\varepsilon_{n}<\frac{\kappa^{2}}{L}\sqrt{(1-\cos(\pi/12))\cos(\pi/12)}$ (98)
Therefore, using Lemma B.1, there must exist $i_{1}^{\prime},\dots,i_{d}^{\prime}\in[n]$ such that $\|\mathbf{u}_{i_{k}^{\prime}}-\mathbf{z}_{k}\|\leq\varepsilon_{n}$ for all $k\in[d]$, which in turn ensures existence of $i_{1},\dots,i_{d}\in\mathcal{U}^{\star}$ such that $\|\mathbf{u}_{i_{k}}-\mathbf{z}_{k}\|\leq 7\varepsilon_{n}$. Finally, using property (i) together with triangular inequality ensures (97). Existence of $j_{1},\dots,j_{k}\in\mathcal{V}^{\star}$ satisfying (97) is proven similarly.
Consequently, we obtain that for all $k,\ell\in[d]$,
$\|\mathbf{u}_{i}-\mathbf{v}_{j_{\ell}}\|_{2}\leq\sqrt{2}+8\varepsilon_{n},\quad\|\mathbf{u}_{i_{k}}-\mathbf{v}_{j_{\ell}}\|_{2}\leq\sqrt{2}+16\varepsilon_{n},\quad\|\mathbf{u}_{i_{k}}-\mathbf{v}_{j}\|_{2}\leq\sqrt{2}+8\varepsilon_{n}$ (99)
and by Lemma E.5 together with property $(iii)$, we obtain
$\sigma_{d}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)=\sigma_{d}\left(\left(\theta(\mathbf{u}_{i_{k}})^{\top}\phi(\mathbf{v}_{j_{\ell}})\right)_{k,\ell\in[d]}\right)\geq\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{4}$ (100)
provided the following condition holds
$8\varepsilon_{n}<\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{8BL}$ (101)
Since $\Sigma_{n}$ is an $\varepsilon_{n}$-net, using using (99), there exists $\mathbf{z}_{i},\mathbf{z}_{i_{1}},\dots,\mathbf{z}_{i_{d}}$ and $\mathbf{z}_{j},\mathbf{z}_{j_{1}},\dots,\mathbf{z}_{j_{d}}$ in $\Sigma_{n}$ such that for all $k,\ell\in[d]$, we have
$\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i}),\quad\mathbf{u}_{i_{k}}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i_{k}}),\quad\mathbf{v}_{j}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j}),\quad\mathbf{v}_{i_{\ell}}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i_{\ell}})$ (102)
and
$\|\mathbf{z}_{i}-\mathbf{z}_{j_{\ell}}\|_{2}\leq\sqrt{2}+10\varepsilon_{n},\quad\|\mathbf{z}_{i_{k}}-\mathbf{z}_{j_{\ell}}\|_{2}\leq\sqrt{2}+18\varepsilon_{n},\quad\|\mathbf{z}_{i_{k}}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+10\varepsilon_{n},$ (103)
thus, given the choice of $N_{n}$, we may apply Lemma B.2 to conclude that the event:
$\forall i^{\prime}\in\mathcal{I},\forall j^{\prime}\in\mathcal{J},\qquad\bar{h}_{ij^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j}\neq\star,$
holds with probability at least $1-\delta$, provided that $\varepsilon_{n}<(\sqrt{3}-\sqrt{2})/20$ and
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\geq\frac{C}{n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right)$ (104)
This concludes the proof. ∎
Imputation.
We are now ready to present Lemma D.2 where we establish finite-sample guarantees on our matrix imputation procedure.
Lemma D.2.
Suppose that $\theta(\cdot)$, ϕ()italic-$\phi(\cdot)$, and $N_{n}$ satisfy the same assumptions as in Lemma D.1. Let $\delta\in(0,1)$. Then, the following event:
$\max_{(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}}|\hat{h}_{ij}-h_{ij}|\leq c\left(\frac{d^{3}B^{2}}{(m^{2}\kappa^{4}\wedge m^{4}\kappa^{8})}\wedge 1\right)\left(LB\varepsilon_{n}+\frac{\sigma}{n}\left(\frac{2}{\varepsilon_{n}}\right)^{d-1}\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}\right)$
holds with probability at least $1-\delta$, provided that
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}$ $\geq\frac{Cd}{m^{2}\kappa^{4}n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right)$ (105)
$\varepsilon_{n}$ $\geq\frac{C}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)}$ (106)
$\varepsilon_{n}$ $\leq C\left(\frac{\kappa^{4}(m^{2}\wedge 1)}{dL(B\vee 1)}\wedge 1\right),$ (107)
$n$ $\geq Cd\log\left(\frac{e\,n}{\delta}\right),$ (108)
for some universal constants $c,C>0$.
Proof of Lemma D.2.
Fix $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ such that $\hat{h}_{ij}=\star$.
Step 1. (Revealed entries of interest). By Lemma D.1, the event: there exists $\mathcal{I}\subseteq\mathcal{U}^{\star}$, $\mathcal{J}\subseteq\mathcal{V}^{\star}$ such that $|\mathcal{I}|=|\mathcal{J}|=d$, and
$\forall i^{\prime}\in\mathcal{I},\forall j^{\prime}\in\mathcal{J},\qquad\bar{h}_{ij^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j}\neq\star,\qquad\text{and}\qquad\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})\geq c\,m^{2}\kappa^{4}$ (109)
holds with probability at least $1-\delta$, where $c>0$ is an absolute constant, provided $\varepsilon_{n}$ satisfies the condition (96).
Step 2. (Error bound for revealed entries). Next, we know by Lemma C.1, that the event: for all entries $(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$, such that $\bar{h}_{ij}\neq\star$,
$|\bar{h}_{ij}-h_{ij}|\leq\delta_{\operatorname{noise}}(n)=14LB\varepsilon_{n}+\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\log\left(\frac{2n^{2}}{\delta}\right)}$ (110)
holds with probability at least $1-\delta$, provided that $\varepsilon_{n}$ satisfies the condition (89) and $n$ is large enough as described in (90).
Step 3. (Error bound for imputed entries). Now, we are ready to use Lemma D.3, which, under the intersection of the two events described by and , entails that
$|\hat{h}_{ij}-h_{ij}|$ $\leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\delta_{\operatorname{noise}}(n)$ (111)
$\leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{cm^{2}\kappa^{4}}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{c^{2}m^{4}\kappa^{8}}\right)\delta_{\operatorname{noise}}(n)$ (112)
provided that the condition $\delta_{\operatorname{noise}}(n)d\leq cm^{2}\kappa^{4}/4$ holds, which we may be enforced by assuming that
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\geq\frac{Cd}{m^{2}\kappa^{4}n}\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}\qquad\text{and}\qquad\varepsilon_{n}\leq\frac{Cm^{2}\kappa^{4}}{dLB}$ (113)
for some universal constant $C>0$ sufficiently large. ∎
Perturbation bound.
The following Lemma is adapted from Proposition 13 of Shah et al. [2020].
Lemma D.3.
Let $\mathbf{H}$ be an $|\mathcal{U}^{\star}|\times|\mathcal{V}^{\star}|$ dimensional matrix with entries are in $[-dB^{2},dB^{2}]$, and satisfying $\mathsf{rank}(\mathbf{H})=d$. Let $\overline{\mathbf{H}}$ be an alternative $|\mathcal{U}^{\star}|\times|\mathcal{V}^{\star}|$ dimensional matrix. Now, for any $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$, assume that there exists $\mathcal{I}\subseteq\mathcal{U}^{\star}$ and $\mathcal{J}\subseteq\mathcal{V}^{\star}$, such that $|\mathcal{I}|=|\mathcal{J}|=d$, and for all $k\in\mathcal{I},\ell\in\mathcal{J}$,
$|\mathbf{H}_{i\ell}-\overline{\mathbf{H}}_{i\ell}|\leq\epsilon,\qquad|\mathbf{H}_{k\ell}-\overline{\mathbf{H}}_{k\ell}|\leq\epsilon,\qquad|\mathbf{H}_{kj}-\overline{\mathbf{H}}_{kj}|\leq\epsilon,$ (114)
for some $0<\epsilon\leq\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})/4d$. Then, defining
$\widehat{\mathbf{H}}_{ij}=\overline{\mathbf{H}}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\overline{\mathbf{H}}_{\mathcal{I},j},$ (115)
it holds that
$\left\lvert\widehat{\mathbf{H}}_{i,j}-\mathbf{H}_{i,j}\right\rvert$ $\leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\epsilon$ (116)
Proof.
First, observe that for any $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$, we have by Lemma H.8 that
$\mathbf{H}_{ij}=\mathbf{H}_{i,\mathcal{J}}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}$ (117)
as $\mathbf{H}_{\mathcal{I},\mathcal{J}}$ and $\mathbf{H}$ both have rank $d$.
Next, we fix some $i\in\mathcal{U}^{\star},v\in\mathcal{V}^{\star}$ and consider the error $\widehat{\mathbf{H}}_{ij}-\mathbf{H}_{ij}$. According to the definition of $\widehat{\mathbf{H}}$ and (117),
$\widehat{\mathbf{H}}_{i,j}-\mathbf{H}_{i,j}=\overline{\mathbf{H}}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\overline{\mathbf{H}}_{\mathcal{I},j}-\mathbf{H}_{i,\mathcal{J}}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}$ (118)
$\quad\quad=\overline{\mathbf{H}}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\overline{\mathbf{H}}_{\mathcal{I},j}-\mathbf{H}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}+\mathbf{H}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}-\mathbf{H}_{i,\mathcal{J}}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}$ (119)
$\quad\quad={\sf Tr}\left(\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\left(\overline{\mathbf{H}}_{\mathcal{I},j}\overline{\mathbf{H}}_{i,\mathcal{J}}-\mathbf{H}_{\mathcal{I},j}\mathbf{H}_{i,\mathcal{J}}\right)\right)+{\sf Tr}\left(\left((\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right)\left(\mathbf{H}_{\mathcal{I},j}\mathbf{H}_{i,\mathcal{J}}\right)\right)$ (120)
Since $\lvert{\sf Tr}(AB)\rvert\leq\sqrt{\mathsf{rank}(B)}\left\lVert A\right\rVert_{\operatorname{op}}\left\lVert B\right\rVert_{F}$, we obtain
$\left\lvert\widehat{\mathbf{H}}_{i,j}-\mathbf{H}_{i,j}\right\rvert$ $\leq\sqrt{2}\left\lVert\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\right\rVert_{\operatorname{op}}\left\lVert\overline{\mathbf{H}}_{\mathcal{I},j}\overline{\mathbf{H}}_{i,\mathcal{J}}-{\mathbf{H}}_{\mathcal{I},j}{\mathbf{H}}_{i,\mathcal{J}}\right\rVert_{F}$ (121)
$+\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}\left\lVert\mathbf{H}_{\mathcal{I},j}\mathbf{H}_{i,\mathcal{J}}\right\rVert_{F}.$ (122)
Moving forward, we will separately establish upper bounds for the two terms. To upper bound (121), note first that by assumption $\left\lVert\Delta\right\rVert_{\operatorname{op}}\leq 2d\left\lVert\Delta\right\rVert_{\operatorname{max}}\leq 2\epsilon d$, where $\Delta=\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}-\mathbf{H}_{\mathcal{I},\mathcal{J}}$. Since $\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})\geq\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})-\left\lVert\Delta\right\rVert_{\operatorname{op}}$ by Weyl’s inequality, we have
$\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}=\frac{1}{\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})}\leq\frac{1}{\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})-2\epsilon d}\leq\frac{2}{\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})}.$ (123)
To bound the second part of the expression, note $\left\lVert\overline{\mathbf{H}}_{\mathcal{I},j}\overline{\mathbf{H}}_{i,\mathcal{J}}-{\mathbf{H}}_{\mathcal{I},j}{\mathbf{H}}_{i,\mathcal{J}}\right\rVert_{F}\leq 2d(2dB^{2}\epsilon+\epsilon^{2})$ because $\left\lVert\overline{\mathbf{H}}_{i^{\prime},j}\overline{\mathbf{H}}_{i,j^{\prime}}-{\mathbf{H}}_{i^{\prime},j}{\mathbf{H}}_{i,j^{\prime}}\right\rVert_{F}\leq 2dB^{2}\epsilon+\epsilon^{2}$ for all $(i^{\prime},j^{\prime})\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$. Now, to upper bound (122), we first upper bound $\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}$ by using Lemma H.9 to note
$\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}\leq\frac{1+\sqrt{5}}{2}\cdot\frac{4}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\cdot 2\epsilon d.$ (124)
Also it is easy to see that $\left\lVert\mathbf{H}_{\mathcal{I},j}\mathbf{H}_{i,\mathcal{J}}\right\rVert_{F}\leq 2d^{2}B^{2}$. Combining all these bounds together with the condition that $\epsilon d\leq\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})/4$gives
$\left\lvert\widehat{\mathbf{H}}_{i,j}-\mathbf{H}_{i,j}\right\rvert$ $\leq\left(\frac{4\sqrt{2}(d^{2}B^{2}+d\epsilon)}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\epsilon$ (125)
$\leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\epsilon$ (126)
Appendix E Patterns of observations and low-rank structure
In this appendix, we present important structural properties that are verified by the outcome matrix under our setting. More specifically, for any given pair of user-item vectors $\mathbf{u},\mathbf{v}\in\mathcal{S}^{d-1}$, there exits $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{d},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$ such that the following holds:
  • (a)
    The vectors $\mathbf{u},\mathbf{x}_{1},\dots,\mathbf{x}_{d},\mathbf{v},\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ are well separated.
  • (b)
    The matrix $\left(\theta(\mathbf{x}_{i})^{\top}\phi(\mathbf{y}_{j})\right)_{i,j\in[d]}$ is well conditioned.
We made these properties precise in Proposition E.1. Intuitively, property $(a)$ entails that an user-item pair has corresponding latent features $\mathbf{x}_{i},\mathbf{u}_{j}$, then user-item outcome is revealed with high probability, and property $(b)$ entails that the outcome of an user-item with corresponding latent feature $\mathbf{u},\mathbf{v}$ can be well predicted.
E.1 Properties of $\theta$ and $\phi$
We present a couple of properties of $\theta$ and $\phi$ that are key in establishing our result.
Lemma E.1.
If $\theta(\cdot)$ is non-degenerate with parameter $m,\kappa>0$, then the following properties hold:
  • (i)
    The map $\theta$ is an homeomorphism from $\mathcal{S}^{d-1}$ to $\theta(\mathcal{S}^{d-1})$. In fact, it is an $(L,m\kappa)$-bilipschitz map.
  • (ii)
    The map $\tilde{\theta}$ is an homeomorphism from $\mathcal{S}^{d-1}$ to $\mathcal{S}^{d-1}$. In fact, it is an $(L/m,\kappa)$-bilipchitz map.
Proof of Lemma E.1.
To prove (i) and (ii), it suffices to establish that the mappings $\theta$ and $\tilde{\theta}$ are bilipchitz with the appropriate constants which in turn ensures the weaker statement that both mappings are homeomorphisms.
We start by showing (i). Let $\mathbf{x},\mathbf{y}\in\mathcal{S}^{d-1}$, and observe that we have the following elementary identity:
$\|\theta(\mathbf{x})-\theta(\mathbf{y})\|^{2}=\left|\|\theta(\mathbf{x})\|-\|\theta(\mathbf{y})\|\right|^{2}+\|\theta(\mathbf{x})\|\|\theta(\mathbf{y})\|\left\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{y})\right\|^{2}.$ (127)
Given that $\theta(\cdot)$ is non-degenerate with parameters $m,\kappa>0$, it holds that
$\|\theta(\mathbf{x})-\theta(\mathbf{y})\|^{2}\geq\|\theta(\mathbf{x})\|\|\theta(\mathbf{y})\|\left\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{y})\right\|^{2}\geq m^{2}\kappa^{2}\|\mathbf{x}-\mathbf{y}\|^{2}.$ (128)
Thus, it must hold that $\theta$ is $(L,m\kappa)$-bilipschitz. Consequently, $\theta$ is injective and continuous and its inverse is also continuous. Thus it is an homeomorphism.
Now, we prove (ii). First, note that the non-degeneracy property of $\theta$ ensures that $\tilde{\theta}$ is well defined. In addition to that, and using (127) we have for all $\mathbf{x},\mathbf{y}\in\mathcal{S}^{d-1}$,
$L^{2}\|\mathbf{x}-\mathbf{y}\|^{2}\geq\|\theta(\mathbf{x})-\theta(\mathbf{y})\|^{2}\geq m^{2}\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{y})\|^{2},$ (129)
Thus, it follows that $\tilde{\theta}$ is $(L/m,\kappa)$-bilipchitz. This means that $\tilde{\theta}$ is continuous on $\mathcal{S}^{d-1}$, is bijection from $\mathcal{S}^{d-1}$ to $\tilde{\theta}(\mathcal{S}^{d-1})\subseteq\mathcal{S}^{d-1}$, and its inverse is also continuous. This means that $\tilde{\theta}$ is a homeomorphsim and it must be the case that $\tilde{\theta}(\mathcal{S}^{d-1})=\mathcal{S}^{d-1}$. ∎
Lemma E.2.
Let $\mathcal{C}$ be a spherical cap on the unit sphere $\mathcal{S}^{d-1}$ defined by the point $\mathbf{x}$, and polar angle $\alpha$. Then, $\tilde{\theta}(\mathcal{C})$ contains the spherical cap defined by the point $\tilde{\theta}(\mathbf{x})$, and polar angle at least $\tilde{\alpha}=\arccos(1-\kappa^{2}+\kappa^{2}\cos(\alpha))$.
Proof of Lemma E.2.
First, let us define $\tilde{\mathcal{C}}\subseteq\tilde{\theta}(\mathcal{C})$, to be a spherical determined by the point $\tilde{\theta}(\mathbf{x})$ and maximal polar angle $\tilde{\alpha}$ (by maximal we mean that any spherical cap that is defined by the point $\tilde{\theta}(\mathbf{x})$ and has polar angle $\tilde{\alpha}+\varepsilon$ is no more included in $\tilde{\theta}(\mathcal{C})$ for any $\varepsilon>0$).
We claim that $\tilde{\alpha}\geq\arccos(1-\kappa^{2}+\kappa^{2}\cos(\alpha))$. Indeed, since $\tilde{\mathcal{C}}$ has maximal polar angle it must be the case that the boundary of $\tilde{\mathcal{C}}$, denoted $\partial\tilde{\mathcal{C}}$, intersects the boundary $\partial(\tilde{\theta}(\mathcal{C}))$ at least at a point $z$. Moreover, since $\tilde{\theta}$ is a homeomorphism we have $\partial\tilde{\theta}(\mathcal{C})=\tilde{\theta}(\partial\mathcal{C})$, thus $\mathbf{z}\in\partial\tilde{\mathcal{C}}\cap\tilde{\theta}(\partial\mathcal{C})$ and there must exist $\mathbf{x}_{z}\in\partial\mathcal{C}$ such that $\tilde{\theta}(\mathbf{x}_{\mathbf{z}})=\mathbf{z}$. Therefore, we have
$2-2\cos(\tilde{\alpha})=\|\tilde{\theta}(\mathbf{x})-\mathbf{z}\|^{2}=\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{x}_{\mathbf{z}})\|^{2}\geq\kappa^{2}\|\mathbf{x}-\mathbf{x}_{\mathbf{z}}\|^{2}=\kappa^{2}(2-2\cos(\alpha)).$
Rearranging the terms in the above inequality we obtain
$\cos(\tilde{\alpha})\leq(1-\kappa^{2})+\kappa^{2}\cos(\alpha),$
which leads to $\tilde{\alpha}\geq\arccos(1-\kappa^{2}+\kappa^{2}\cos(\alpha))$ (because $\cos(\cdot)$ is a strictly decreasing function from $[0,\pi]$ to $[-1,1]$, and $\alpha,\tilde{\alpha}\in[0,\pi]$). ∎
E.2 Properties of high dimensional simplices
The following result on regular simplicies is borrowed from Parks and Wills [2002] (see their Theorem 1).
Lemma E.3.
Let $\mathbf{a}_{1},\dots,\mathbf{a}_{d}$ be unit vectors such that $\{\mathbf{a}_{1},\dots,\mathbf{a}_{d}\}$ form an $(d-1)$-regular simplex embedded in $\mathbb{R}^{d}$. Then, we have for all $i,j\in[d]$ such that $i\neq j$, we have
$\langle\mathbf{a}_{i},\mathbf{a}_{j}\rangle=\frac{-1}{d-1}.$
Below, we present a result on the singular values of a matrix formed by a set of linearly independent vectors that are equidistant.
Lemma E.4.
Let $\mathbf{x}_{1},\dots,\mathbf{x}_{d}\in\mathcal{S}^{d-1}$ be unit vectors that are equidistant, i.e. there exists a constant $c>0$, such that for all $i,j\in[d]$, $i\neq j$, $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|=c$, then, denoting $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\cdots&\mathbf{x}_{d}\end{bmatrix}$, we have
[…]
Proof of Lemma E.4.
Let us start by observing that, for all $i,j\in[d]$,
$(\mathbf{X}^{\top}\mathbf{X})_{ij}=\begin{cases}1&\text{ if }i=j,\\ (2-c^{2})/2&\text{ if }i\neq j.\end{cases}$ (130)
where we used the fact that $(\mathbf{X}^{\top}\mathbf{X})_{ij}=\mathbf{x}_{i}^{\top}\mathbf{x}_{j}=(\|\mathbf{x}_{i}\|^{2}+\|\mathbf{x}_{j}\|^{2}-\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2})/2$. Next, we will provide $\mathbf{z}_{1},\dots,\mathbf{z}_{d}$, $d$ orthogonal eigenvectors of $\mathbf{X}^{\top}\mathbf{X}$, which will allow us to explicitely provide the singular values of $\mathbf{X}$. We define
$\mathbf{z}_{1}=\begin{bmatrix}1\\ 1\\ 1\\ \vdots\\ 1\end{bmatrix},\mathbf{z}_{2}=\begin{bmatrix}-1\\ d-1\\ -1\\ \vdots\\ -1\end{bmatrix},\mathbf{z}_{3}=\begin{bmatrix}-1\\ -1\\ d-1\\ \vdots\\ -1\end{bmatrix},\dots,\mathbf{z}_{d}=\begin{bmatrix}-1\\ -1\\ -1\\ \vdots\\ d-1\end{bmatrix}$
It is not difficult to verify that $\mathbf{z}_{1},\dots,\mathbf{z}_{d}$ are linearly independent. Furthermore, we can also verify that
$(\mathbf{X}^{\top}\mathbf{X})\mathbf{z}_{1}=(1+(d-1)(2-c^{2})/2)\mathbf{z}_{1},$
and for all $i\in\{2,\dots,d\}$, we have
$(\mathbf{X}^{\top}\mathbf{X})\mathbf{z}_{i}=(c^{2}/2)\mathbf{z}_{i}.$
This proves that $\mathbf{X}^{\top}\mathbf{X}$ has two eigenvalues such that the smallest one $(c^{2}/2)$ has multiplicity $d-1$. We conclude that
$\sigma_{1}(\mathbf{X})=\sqrt{(1+(d-1)(2-c^{2})/2)},\quad\text{and}\quad\sigma_{2}(\mathbf{X})=\cdots=\sigma_{d}(\mathbf{X})=c/\sqrt{2}$
E.3 Existence of a well conditioned sub-matrix of size $d\times d$
Below, we describe the algorithm to find $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ with the desired properties highlighted in Proposition E.1.
Input: $\mathbf{u},\mathbf{v}\in\mathcal{S}^{d-1}$
1 Step 1: Finding sufficiently close features
2 Find the geodesic $\Gamma$ between $\mathbf{u}$ and $\mathbf{v}$ on the hypersphere $\mathcal{S}^{d-1}$;
3 Find the midpoint $\mathbf{m}\in\mathcal{S}^{d-1}$ of $\Gamma$ ;
4 Find the point $\mathbf{x}$ on $\Gamma$ such that $\arccos(\langle\mathbf{x},\mathbf{m}\rangle)=\pi/8$ and is closer to $\mathbf{v}$ ;
5 Find the point $\mathbf{y}$ on $\Gamma$ such that $\arccos(\langle\mathbf{y},\mathbf{m}\rangle)=\pi/8$ and is closer to $\mathbf{u}$ ;
6 Let $\mathcal{C}(\mathbf{x},\pi/12)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\mathbf{x}$ and polar angle $\pi/12$;
7 Let $\mathcal{C}(\mathbf{y},\pi/12)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\mathbf{y}$ and polar angle $\pi/12$;
8 Step 2: Finding well separated features
9 Let $\alpha=\arccos(1-\kappa^{2}+\kappa^{2}\cos(\pi/12))$;
10 Let $\mathcal{C}(\tilde{\theta}(\mathbf{x}),\alpha)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\tilde{\theta}(\mathbf{x})$ and polar angle $\alpha$;
11 Let $\mathcal{C}(\tilde{\phi}(\mathbf{y}),\alpha)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\tilde{\phi}(\mathbf{y})$ and polar angle $\alpha$;
12 Let $\mathbf{z}_{1},\dots,\mathbf{z}_{d}$ be vertices of a regular $(d-1)$-simplex inscribed in $\partial\mathcal{C}(\tilde{\theta}(\mathbf{x}),\alpha)$;
13 Let $\mathbf{w}_{1},\dots,\mathbf{w}_{d}$ be vertices of a regular $(d-1)$-simplex inscribed in $\partial\mathcal{C}(\tilde{\phi}(\mathbf{y}),\alpha)$;
14 For all $i\in[d]$, set $\mathbf{x}_{i}=\tilde{\theta}^{-1}(\mathbf{z}_{i})$ and $\mathbf{y}_{i}=\tilde{\phi}^{-1}(\mathbf{w}_{i})$;
Output: $\mathbf{x}_{1},\dots,\mathbf{x}_{d},\mathbf{y}_{1},\dots,\mathbf{y}_{d}\in\mathcal{S}^{d-1}$
Algorithm 1 Search rank $d$-sub-matrix in $\mathbf{X}$
Proposition E.1.
For all $\mathbf{u},\mathbf{v}\in\mathcal{S}^{d-1}$, there exists $\mathbf{x}_{1},\dots,\mathbf{x}_{d},\mathbf{y}_{1},\dots,\mathbf{y}_{d}\in\mathcal{S}^{d-1}$, such that:
  • (i)
    for all $i,j\in[d]$, $\|\mathbf{v}-\mathbf{x}_{i}\|_{2}\leq\sqrt{2}$, $\|\mathbf{u}-\mathbf{y}_{j}\|_{2}\leq\sqrt{2}$ and $\|\mathbf{x}_{i}-\mathbf{y}_{j}\|_{2}\leq\sqrt{2}$,
  • (ii)
    for all $i,j\in[d]$, $i\neq j$, we have $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}\wedge\|\mathbf{y}_{i}-\mathbf{y}_{j}\|_{2}\geq L^{-1}\kappa^{2}\sqrt{(1-\cos(\pi/12))\cos(\pi/12)}$,
  • (iii)
    and the following holds
    $\sigma_{d}\left(\begin{bmatrix}\theta(\mathbf{x}_{1})&\dots&\theta(\mathbf{x}_{d})\end{bmatrix}^{\top}\begin{bmatrix}\phi(\mathbf{y}_{1})&\ldots&\phi(\mathbf{y}_{d})\end{bmatrix}\right)\geq\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{2}$
Proof of Proposition E.1.
We prove existence of $\mathbf{x}_{1},\dots,\mathbf{x}_{d},\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ satisfying properties $(i)$ and $(ii)$ by construction. This construction is geometric in nature and draws inspiration from properties of high dimensional simplices. We provide an algorithmic description of such construction in Algorithm 1.
Claim 1. Step 1 in Algorithm 1, ensures that any choice of points $\mathbf{x}_{1},\dots,\mathbf{x}_{d}\in\mathcal{C}(\mathbf{x},\pi/12)$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}\in\mathcal{C}(\mathbf{y},\pi/12)$ satisfies property (i).
Proof of Claim 1. Let $\mathbf{x}^{\prime}\in\mathcal{C}(\mathbf{x},\pi/12)$ and $\mathbf{y}^{\prime}\in\mathcal{C}(\mathbf{y},\pi/12)$. Note that by construction $\langle\mathbf{x},\mathbf{v}\rangle\geq\cos(\pi/2-\pi/8)$ (because $\mathbf{m},\mathbf{x},\mathbf{v}$ are coplanar and $\langle\mathbf{x},\mathbf{m}\rangle\geq\cos(\pi/2)$). Let $\mathbf{x}^{\prime}\in\mathcal{C}(\mathbf{x},\pi/12)$. Then, we have $\langle\mathbf{v},\mathbf{x}^{\prime}\rangle\geq\cos(3\pi/8+\pi/12)\geq 0$. This entails that $\|\mathbf{x}^{\prime}-\mathbf{v}\|\leq\sqrt{2}$. We can prove similarly that for all $\mathbf{y}^{\prime}\in\mathcal{C}(\mathbf{y},\pi/12)$, we have $\|\mathbf{y}^{\prime}-\mathbf{u}\|\leq\sqrt{2}$. Next, note also that $\langle\mathbf{x}^{\prime},\mathbf{y}^{\prime}\rangle\geq\cos(2\pi/12+2\pi/8)\geq 0$, thus $\|\mathbf{x}^{\prime}-\mathbf{y}^{\prime}\|\leq\sqrt{2}$. This concludes the proof of the claim.
Claim 2. Step 2 in Algorithm 1, ensures that the selected vectors $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ satisfy property (iii).
Proof of Claim 3. First, we note that Lemma E.1 and Lemma E.2 ensure that $\tilde{\theta}$ and $\tilde{\phi}$ are well defined, and that $\mathcal{C}(\tilde{\theta}(\mathbf{x}),\alpha)\subseteq\tilde{\theta}(\mathcal{C}(\mathbf{x},\pi/12))$ and $\mathcal{C}(\tilde{\phi}(\mathbf{y}),\alpha)\subseteq\tilde{\phi}(\mathcal{C}(\mathbf{y},\pi/12))$. This in turn implies that $\mathbf{x}_{1},\dots,\mathbf{x}_{d}\in\mathcal{C}(\mathbf{x},\pi/12)$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}\in\mathcal{C}(\mathbf{y},\pi/12)$. It remains to lower bound the the smallest singular value of $\begin{bmatrix}\theta(\mathbf{x}_{1})&\ldots&\theta(\mathbf{x}_{d})\end{bmatrix}^{\top}\begin{bmatrix}\phi(\mathbf{y}_{1})&\ldots&\phi(\mathbf{y}_{d})\end{bmatrix}$. Under assumption A2, we have
$\sigma_{d}\left(\begin{bmatrix}\theta(\mathbf{x}_{1})&\ldots&\theta(\mathbf{x}_{d})\end{bmatrix}\right)^{2}$ $=\inf_{\mathbf{z}\in\mathcal{S}^{d-1}}\sum_{i=1}^{d}|\theta(\mathbf{x}_{i})^{\top}\mathbf{z}|^{2}$
$\geq m^{2}\inf_{\mathbf{z}\in\mathcal{S}^{d-1}}\sum_{i=1}^{d}|\tilde{\theta}(\mathbf{x}_{i})^{\top}\mathbf{z}|^{2}$
$\geq m^{2}\sigma_{d}\left(\begin{bmatrix}\tilde{\theta}(\mathbf{x}_{1})&\ldots&\tilde{\theta}(\mathbf{x}_{d})\end{bmatrix}\right)^{2}$
where we used the variational definition of the smallest singular value. Now, observe that by construction $\tilde{\theta}(\mathbf{x}_{1}),\dots,\tilde{\theta}(\mathbf{x}_{d})$ are the vertices of an $(d-1)$-simplex inscribed in $\partial\mathcal{C}(\tilde{\phi},\alpha)$. Thus, for each $i\neq j$, we have
$\left\|\tilde{\theta}(\mathbf{x}_{i})-\tilde{\theta}(\mathbf{x}_{j})\right\|=\sqrt{1-\cos^{2}(\alpha)}.$
Therefore, we may apply Lemma E.4 and obtain:
$\sigma_{d}\left(\begin{bmatrix}\tilde{\theta}(\mathbf{x}_{1})&\ldots&\tilde{\theta}(\mathbf{x}_{d})\end{bmatrix}\right)^{2}\geq\frac{1-\cos^{2}(\alpha)}{2}$
We can similarly lower bound $\sigma_{d}\left(\begin{bmatrix}\tilde{\phi}(\mathbf{y}_{1})&\ldots&\tilde{\phi}(\mathbf{y}_{d})\end{bmatrix}\right)$. We conclude that
$\sigma_{d}\left(\begin{bmatrix}\theta(\mathbf{x}_{1})&\ldots&\theta(\mathbf{x}_{d})\end{bmatrix}^{\top}\begin{bmatrix}\phi(\mathbf{y}_{1})&\ldots&\phi(\mathbf{y}_{d})\end{bmatrix}\right)$ $\geq\frac{m^{2}(1-\cos^{2}(\alpha))}{\sqrt{2}}$
$=\frac{m^{2}\left(1-\left(1-\kappa^{2}+\kappa^{2}\cos(\pi/12)\right)^{2}\right)}{\sqrt{2}}$
$\geq\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{\sqrt{2}}.$
Claim 3. Step 2 in Algorithm 1 also ensures that the selected vectors $\mathbf{x}_{1},\dots,\mathbf{x}_{d},\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ satisfy property $(ii)$.
Indeed, for all $i\neq j$, we have by construction that $\|\tilde{\theta}(\mathbf{x}_{i})-\tilde{\theta}(\mathbf{x}_{j})\|_{2}=\sqrt{1-\cos^{2}(\alpha)}$. Then, using the fact that $\widetilde{\theta}(\cdot)$ is $L$-lipschitz which comes from Lemma E.1, we can immediately lower bound $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}$ as claimed. The claim for $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ follows similarly.
Proposition E.1 guarantees existence of $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$ such that
$\sigma_{d}\left(\begin{bmatrix}\theta(\mathbf{x}_{1})&\ldots&\theta(\mathbf{x}_{d})\end{bmatrix}\right)\wedge\sigma_{d}\left(\begin{bmatrix}\phi(\mathbf{y}_{1})&\dots&\phi(\mathbf{y}_{d})\end{bmatrix}\right)\geq C$
for some universal positive constant $C>0$. However, we may only have approximate values of the vectors $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$. Below, we present a result showing that this does not matter too much.
Lemma E.5.
Let $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$ such that
$\sigma_{d}\left(\theta(\mathbf{X})^{\top}\phi(\mathbf{Y})\right)\geq C.$
Let $\mathbf{u}_{1},\dots,\mathbf{u}_{d}$ and $\mathbf{v}_{1},\dots,\mathbf{v}_{d}\in\mathcal{S}^{d-1}$ be such that
[…]
for some $\varepsilon<C/(4BL)$. Then, we have
$\sigma_{d}\left(\theta(\mathbf{U})^{\top}\phi(\mathbf{V})\right)\geq\frac{C}{2}$
Proof of Lemma E.5.
As $\theta$ and $\phi$ are $L$-Lipschitz, we know
$\left\lvert\theta(\mathbf{u}_{i})^{\top}\phi(\mathbf{v}_{j})-\theta(\mathbf{x}_{i})^{\top}\phi(\mathbf{y}_{j})\right\rvert\leq 2BL\varepsilon$
Thus, applying Weyl’s inequality, we have
$\left\lvert\sigma_{d}(\theta(\mathbf{U})^{\top}\phi(\mathbf{V}))-\sigma_{d}(\theta(\bm{X})^{\top}\phi(\bm{Y}))\right\rvert$ $\leq\sigma_{1}(\sigma_{d}(\theta(\mathbf{U})^{\top}\phi(\mathbf{V}))-\theta(\mathbf{X})^{\top}\phi(\mathbf{Y}))$
$\leq\|\theta(\mathbf{U}_{n})^{\top}\phi(\mathbf{V}_{n})-\theta(\mathbf{X})^{\top}\phi(\mathbf{Y})\|_{\max}$
$\leq\max_{i,j\in[d]}\left\lvert\theta(\mathbf{u}_{i})^{\top}\phi(\mathbf{v}_{j})-\theta(\mathbf{x}_{i})^{\top}\phi(\mathbf{y}_{j})\right\rvert$
$\leq 2BL\varepsilon$
If $\varepsilon\leq C/(4BL)$, then we have
$\sigma_{d}\left\lparen\theta(\mathbf{U})^{\top}\phi(\mathbf{V})\right\rparen\geq\frac{C}{2}$
Appendix F Proof of Theorem 4.1
Fix $i,j\in[n]$. We know, thanks to our clustering procedure, that there exists clusters $k\in\mathcal{U}^{\star}$, $\ell\in\mathcal{V}^{\star}$ such that $i\in\mathcal{U}_{k}$ and $j\in\mathcal{V}_{\ell}$. Thus, we may decompose the prediction error as follows
$x_{ij}-\hat{x}_{ij}=x_{ij}-h_{k\ell}+h_{k\ell}-\hat{h}_{k\ell}.$ (131)
Step 1. (Bounding the first term). First, we note that
$|h_{k\ell}-x_{ij}|$ $=|f(\mathbf{u}_{k},\mathbf{v}_{\ell})-f(\mathbf{u}_{i},\mathbf{v}_{j})|\leq\sqrt{d}BL(\|\mathbf{u}_{k}-\mathbf{v}_{i}\|_{2}+\|\mathbf{v}_{\ell}-\mathbf{v}_{j}\|_{2}).$ (132)
We know, under our clustering procedure, that $\hat{d}_{ki}^{\mathbf{u}}\leq 6\varepsilon_{n}$ and $\hat{d}_{\ell j}^{\mathbf{v}}\leq 6\varepsilon_{n}$. Hence, using Lemma A.1, we obtain that the event:
$|h_{k\ell}-x_{ij}|\leq\sqrt{d}BL(\|\mathbf{u}_{k}-\mathbf{u}_{i}\|_{2}+\|\mathbf{v}_{\ell}-\mathbf{v}_{j}\|_{2})\leq 14\sqrt{d}BL\varepsilon_{n}$ (133)
holds with probability $1-\delta/(2n^{2})$, provided that provided that
$\varepsilon_{n}$ $\geq\frac{c}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)}=cn^{2\beta-\frac{1}{2}}\sqrt{d^{3}\log\left(\frac{e\,n}{\delta}\right)}$ (134)
$n$ $\geq Cd\log\left(\frac{e\,n}{\delta}\right),$ (135)
where $c,C>0$ are universal constants, and recall that $\rho_{n}=n^{-\beta}$. Note that if we wish to obtain a non-trivial error rate, then condition (134) implies that we require
$\beta<\frac{1}{4}.$
Step 2. (Bounding the second term). Second, we invoke Lemma D.2, to obtain that the event:
$|\hat{h}_{k\ell}-h_{k\ell}|\leq c\left(\frac{d^{3}B^{2}}{(m^{2}\kappa^{4}\wedge m^{4}\kappa^{8})}\wedge 1\right)\left(LB\varepsilon_{n}+\frac{\sigma}{n}\left(\frac{2}{\varepsilon_{n}}\right)^{d-1}\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}\right)$ (136)
holds with probability at least $1-\delta/(2n^{2})$, provided that (134), (135), and
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}$ $\geq\frac{Cd}{m^{2}\kappa^{4}n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right)=\frac{Cd}{m^{2}\kappa^{4}n}\left(n^{\beta/2}\log^{1/2}\left(\frac{en}{\delta}\right)+\log\left(\frac{en}{\delta}\right)\right)$ (137)
$\varepsilon_{n}$ $\leq C\left(\frac{\kappa^{4}(m^{2}\wedge 1)}{dL(B\vee 1)}\wedge 1\right)$ (138)
where $C>0$ is some universal constant.
Step 3. (Putting everything together). To complete our analysis, we choose $\varepsilon_{n}=2n^{-\alpha}$, and find the optimal choice of $\alpha$ that minimizes prediction error. Under the two events described above in step 1 and step 2, ignoring the problem dependent constants, it is clear that the prediction error scales as follows with $n$,
$|\hat{x}_{ij}-x_{ij}|$ $\lesssim\varepsilon_{n}+\frac{1}{n}\left(\frac{2}{\varepsilon_{n}}\right)^{d-1}\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}$ (139)
$\lesssim\frac{1}{n^{\alpha}}+n^{\alpha(d-1)-1+\frac{\beta}{2}}\sqrt{\log\left(\frac{e\,n}{\delta}\right)}.$ (140)
Clearly, choosing
$\alpha=\frac{2-\beta}{2d},$ (141)
yields the minimal error rate
$|\hat{x}_{ij}-x_{ij}|\lesssim n^{-\frac{2-\beta}{2d}}\sqrt{\log\left(\frac{e\,n}{\delta}\right)}.$ (142)
However, for such choice of $\alpha$ to be valid, we need to verify that the conditions (134), (137), and (138) are verified. Now, we can easily verify that condition (134) holds if
$\beta<\frac{d-1}{4d-1}\left(1-\frac{d}{d-1}\frac{\log\left(c^{2}d^{3}\log\left(en/\delta\right)\right)}{\log(n)}\right)$ (143)
Next, we can also easily verify that condition (137) holds if
$\beta$ $\leq 2\left(1-\frac{\log\left(\frac{Cd}{2m^{2}\kappa^{4}}\sqrt{\log\left(\frac{en}{\delta}\right)}\right)}{\log(n)}\right)$ (144)
$\frac{1}{d}$ $\geq\frac{\log(Cd/(2m^{2}\kappa^{4}))+\log\log(en/\delta)}{\log(n)}$ (145)
Finally, we easily verify that condition (138) holds if
$\beta$ $\leq 2\left(1+\frac{d\log\left(\frac{C}{2}\left(\frac{\kappa^{4}(m^{2}\wedge 1)}{dL(B\vee 1)}\wedge 1\right)\right)}{\log(n)}\right).$ (146)
In summary, we see that when
$n\geq\left(\frac{C^{\prime}d^{3}LB}{m^{2}\kappa^{4}}\log\left(\frac{e\,n}{\delta}\right)\right)^{2d},\quad d>9/4,\quad\beta<\frac{1}{4},$ (147)
then all the conditions above are satisfied, and thus (134), (137), and (138) are verified.
Appendix G Psuedocode
Pseudocode for the main algorithm MNN and subroutine Impute are as follows:
Input: $\mathbf{A},\mathbf{Y}$
1 Distance Estimation:
2 For all $i,j\in[n]$ define the distance estimates $\hat{d}^{\mathbf{u}}_{ij}$ and $\hat{d}^{\mathbf{v}}_{ij}$ as described in (3.1);
3 Clustering:
4 Initialize the central users $\mathcal{U}^{\star}$ to be an empty set;
5 While $\hat{d}^{\mathbf{u}}_{ij}>6\varepsilon_{n}$ for all $j\in\mathcal{U}^{\star}$, add $i$ to $\mathcal{U}^{\star}$;
6 For each $i\in\mathcal{U}^{\star}$, initialize empty set $\mathcal{U}_{i}$;
7 For each $i\in\mathcal{U}^{\star}$ and each user $k\in[n]$, if $\hat{d}^{\mathbf{u}}_{ik}\leq\hat{d}^{\mathbf{u}}_{jk}$ for all $j\in\mathcal{U}^{\star}$, add $k$ to $\mathcal{U}_{i}$;
8 Run analogous procedure on items to create central items $\mathcal{V}^{\star}$ and sets $\mathcal{V}_{j},j\in\mathcal{V}^{\star}$;
9 Form partitions $\Omega_{ij}=\{(k,\ell):a_{kl}=1,k\in\mathcal{U}_{i},\ell\in\mathcal{V}_{j}\}$;
10 Matrix Completion:
11 Construct the unfilled clustered outcome matrix $\overline{\mathbf{H}}$ as in (8);
12 Construct the imputed clustered outcome matrix $\widehat{\mathbf{H}}$ as in (9);
Output: $(\hat{x}_{ij})_{i,j\in[n]}$
Algorithm 2 MNN
Input: $\overline{\mathbf{H}}$, $(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$
1 Step 1:
2 Search exhaustively for ×$\mathcal{I}\times\mathcal{J}\subseteq\mathcal{U}^{\star}\times\mathcal{V}^{\star}$ such that $|\mathcal{I}|=|\mathcal{J}|=d$ and $\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})$ is maximized;
3 Step 2:
4 Set $\hat{h}_{ij}=\begin{cases}\overline{\mathbf{H}}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\overline{\mathbf{H}}_{\mathcal{I},j}&\text{if }\mathcal{I},\mathcal{J}\text{ are found}\\ \star&\text{otherwise}\end{cases}$;
Output: $\hat{h}_{ij}$
Algorithm 3 $\texttt{Impute}(\bar{\mathbf{H}},i,j)$
A visual description of the way MNN works is as follows:
$\ldots$
$\ldots$
$\ldots$
$\ldots$
Figure 5: A visual depiction of the way MNN processes the users and items in the outcome matrix. The first matrix depicts the original noisy outcome matrix $\mathbf{Y}$ with several rows and columns highlighted. The second matrix shows how these highlighted rows and columns have been "reordered" based on estimated distances. The final matrix shows the imputed clustered outcome matrix $\hat{\mathbf{H}}$ where all users and items highlighted are assigned to the same central user/item cluster.
Appendix H Helper Lemmas
Lemma H.1.
Let the surface area of the unit $d$-sphere be $A_{d}$. Consider the corresponding hyperspherical cap subtended by the angle $\Phi\in(0,\pi/2)$, and let $A_{d}^{\operatorname{cap}}$ be its surface area. It holds that
$A_{d}^{\operatorname{cap}}/A_{d}\geq\frac{1}{4}\sin^{d-1}\Phi.$ (148)
Proof.
Using a result from Li [2010], we write
$A_{d}^{\operatorname{cap}}/A_{d}=\frac{1}{2}I\left\lparen\sin^{2}\Phi;\frac{d-1}{2},\frac{1}{2}\right\rparen,$ (149)
where $I(x;a,b)$ is known as the regularized incomplete beta function and defined by
$I(x;a,b)=\frac{B(x;a,b)}{B(a,b)}$ (150)
where $B(a,b)$ is the beta function and
[…]
is the incomplete beta function. Then
[…]
$=\frac{2}{d-1}\sin^{d-1}\Phi.$ (154)
Meanwhile, a simple but tight bound Alzer [2001] for the denominator is
$B\left\lparen\frac{d-1}{2},\frac{1}{2}\right\rparen\leq\frac{4}{d-1}.$ (155)
Combining the bounds yields
$A_{d}^{\operatorname{cap}}/A_{d}\geq\frac{1}{4}\sin^{d-1}\Phi.$ (156)
Lemma H.2 (Adapted from Appendix G.4 of Agarwal et al. [2022]).
Let $\mathbf{Y}=\mathbf{M}+\mathbf{E}\in\mathbb{R}^{m\times n}$ with $\mathsf{rank}(\mathbf{M})=d$. Let $\mathbf{U}_{d}\mathbf{\Sigma}_{d}\mathbf{V}_{d}^{\top}$ and $\hat{\mathbf{U}}_{d}\hat{\mathbf{\Sigma}}_{d}\hat{\mathbf{V}}_{d}^{\top}$ denote the top $d$ singular components of the SVD of $\mathbf{M}$ and $\mathbf{Y}$ respectively. Then, the truncated SVD estimator $\hat{\mathbf{M}}=\hat{\mathbf{U}}_{d}\hat{\mathbf{\Sigma}}_{d}\hat{\mathbf{V}}_{d}^{\top}$ is such that for all $j\in[m]$,
$\left\lVert\hat{\mathbf{M}}_{j}-\mathbf{M}_{j}\right\rVert_{2}^{2}\leq 2\cdot\left\lparen\frac{\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}^{2}}{\sigma_{d}^{2}(\mathbf{M})}\left\lparen\left\lVert\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{M}_{j}\right\rVert_{2}^{2}\right\rparen+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen.$ ) . (157)
Proof.
First note that $\hat{\mathbf{M}}=\mathbf{Y}\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}$ by definition of the truncated SVD. Thus we can write
$\hat{\mathbf{M}}_{j}-\mathbf{M}_{j}$ $=(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{Y}_{j}-\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{M}_{j})+(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{M}_{j}-\mathbf{M}_{j})$ (158)
$=\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}+(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{I}_{n})~{}\mathbf{M}_{j}.$ (159)
Since the two terms belong to orthogonal subspaces, we have
$\left\lVert\hat{\mathbf{M}}_{j}-\mathbf{M}_{j}\right\rVert_{2}^{2}=\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{I}_{n})~{}\mathbf{M}_{j}\right\rVert_{2}^{2}.$ (160)
The first term of (160) can be expanded by the triangle inequality as
$\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}$ $\leq 2\cdot\left\lparen\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}-\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen$ ) (161)
$\leq 2\cdot\left\lparen\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\right\rVert_{\operatorname{op}}^{2}\left\lVert\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen.$ ) . (162)
Since $\mathsf{rank}(\mathbf{M})=d$, the Wedin $\sin{\Theta}$ Theorem Wedin [1972] guarantees that
$\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\right\rVert_{\operatorname{op}}\leq\frac{\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}}{\sigma_{d}(\mathbf{M})}.$ (163)
Combining (162) and (163) yields
$\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\leq 2\cdot\left\lparen\frac{\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}^{2}}{\sigma_{d}^{2}(\mathbf{M})}\left\lVert\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen.$ ) . (164)
Now we return to the second term of (160). By definition, $\mathbf{M}=\mathbf{M}\mathbf{V}_{d}\mathbf{V}_{d}^{\top}$. Therefore, using (163) once again, we get
$\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{M}_{j}-\mathbf{M}_{j}\right\rVert_{2}^{2}$ $=\left\lVert(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{V}_{d}\mathbf{V}_{d}^{\top})~{}\mathbf{M}_{j}\right\rVert_{2}^{2}$ (165)
$\leq\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\right\rVert_{\operatorname{op}}^{2}\left\lVert\mathbf{M}_{j}\right\rVert_{2}^{2}$ (166)
$\leq\frac{\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}^{2}}{\sigma_{d}^{2}(\mathbf{M})}\left\lVert\mathbf{M}_{j}\right\rVert_{2}^{2}.$ (167)
Combining (160), (164), and (167) we obtain the desired bound. ∎
Lemma H.3 (Proposition 2.4 of Rudelson and Vershynin [2010]).
Let $\mathbf{X}=[x_{ij}]\in\mathbb{R}^{m\times n}$ be a random matrix whose entries are independent mean-zero sub-gaussian random variables with bounded sub-gaussian norm $\lVert x_{ij}\rVert_{\psi_{2}}\leq 1$. Then for any $t\geq 0$, with probability at least $1-2\exp(-ct^{2})$ it holds that
$\left\lVert\mathbf{X}\right\rVert_{\operatorname{op}}\leq C(\sqrt{m}+\sqrt{n})+t,$ (168)
where $C,c>0$ are absolute constants.
Lemma H.4 (Matrix Bernstein inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be a sequence of independent random symmetric matrices with dimension $d$. Assume that each matrix satisfies
${\rm I}\kern-1.79993pt{\rm E}[\mathbf{X}_{i}]$ $=0,$
$\lambda_{\max}(\mathbf{X}_{i})$ $\leq R.$
Further define
$\sigma^{2}=\left\lVert\sum_{i=1}^{n}{\rm I}\kern-1.79993pt{\rm E}[\mathbf{X}_{i}^{2}]\right\rVert_{\operatorname{op}}.$ (169)
Then
$\mathbb{P}\left[\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{X}_{i}\right)\geq t\right]$ $\leq d\exp{\left(\frac{-t^{2}}{\sigma^{2}+Rt/3}\right)}$
$\leq\begin{cases}d\exp{\left(-t^{2}/4\sigma^{2}\right)}&\quad\text{for $t\leq\sigma^{2}/R$}\\ d\exp{\left(-t^{2}/4R\right)}&\quad\text{for $t\geq\sigma^{2}/R$}\end{cases}.$
Lemma H.5 (Theorem 5.39 of Vershynin [2010]).
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $d<n$, have rows sampled i.i.d. from an isotropic distribution, i.e. ${\rm I}\kern-1.79993pt{\rm E}[\mathbf{X}_{i}\mathbf{X}_{i}^{\top}]=\mathbf{I}_{d}$. Then for every $t\geq 0$, with probability at least $1-2\exp(-ct^{2})$ the singular values of $\mathbf{X}$ satisfy
$\sqrt{n}-C\sqrt{d}-t\leq\sigma_{\min}(\mathbf{X})\leq\sigma_{\max}(\mathbf{X})\leq\sqrt{n}+C\sqrt{d}+t,$
where $C,c>0$ are constants that only depend on the sub-gaussian norm of $\mathbf{X}_{i}$, i.e. $\lVert\mathbf{X}_{i}\rVert_{\psi_{2}}$.
Lemma H.6 (Mean of independent sub-gaussian random variables).
Let $X_{1},\dots,X_{n}$ be a sequence of independent mean-zero random variables with $X_{i}\sim\,\mathsf{subG}\,(\sigma_{i}^{2})$. Then, for any $t>0$, it holds that
$\mathbb{P}\left[\left\lvert\frac{1}{n}\sum_{i=1}^{n}X_{i}\right\rvert>t\right]\leq 2\exp{\left\lparen-\frac{n^{2}t^{2}}{2\sum_{i=1}^{n}\sigma_{i}^{2}}\right\rparen}.$ ) . (170)
Lemma H.7 (Norm of sub-gaussian random vector).
Let $\mathbf{x}\in\mathbb{R}^{n}$ be a mean-zero $\sigma^{2}$-sub-gaussian random vector, i.e. for any fixed $\mathbf{v}\in\mathcal{S}^{n-1}$ the random variable $\mathbf{v}^{\top}\mathbf{x}$ is $\sigma^{2}$-sub-gaussian, or
${\rm I}\kern-1.79993pt{\rm E}[\exp(s\mathbf{v}^{\top}\mathbf{x})]\leq\exp\left\lparen\frac{s^{2}\sigma^{2}}{2}\right\rparen.$ ) . (171)
Then, with probability at least $1-2\exp(-\frac{t^{2}}{2n\sigma^{2}})$, it holds that
$\left\lVert\mathbf{x}\right\rVert_{2}\leq t.$ (172)
Proof.
Since $\mathbf{x}$ is $\sigma^{2}$-sub-gaussian, it is $(n\sigma^{2})$-norm-sub-gaussian as defined Jin et al. [2019]. Using Definition 3 and Lemma 1 of Jin et al. [2019], the result immediately follows. ∎
Lemma H.8 (Lemma 12 of Shah et al. [2020]).
Let $M=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. If $\mathsf{rank}(A)=\mathsf{rank}(M)$, then $D=CA^{\dagger}B$.
Lemma H.9 (Theorem 3.3 of Stewart [1977]).
For any matrices $A,B$ with $B=A+\Delta$, it holds that
$\left\lVert B^{\dagger}-A^{\dagger}\right\rVert_{\operatorname{op}}\leq\frac{1+\sqrt{5}}{2}\max\left(\left\lVert A^{\dagger}\right\rVert_{\operatorname{op}}^{2},\left\lVert B^{\dagger}\right\rVert_{\operatorname{op}}^{2}\right)\left\lVert\Delta\right\rVert_{\operatorname{op}}.$ (173)